### v2

# Polynomial-Time Near-Optimal Estimation over Certain Type-2 Convex Bodies

Abstract
We develop polynomial-time algorithms for near-optimal minimax mean estimation under $\ell_{2}$-squared loss in a Gaussian sequence model under convex constraints. The parameter space is an origin-symmetric, type-2 convex body $K\subset\mathbb{R}^{n}$, and we assume additional regularity conditions: specifically, we assume $K$ is well-balanced, i.e., there exist known radii $r,R>0$ such that $rB_{2}\subseteq K\subseteq RB_{2}$, as well as oracle access to the Minkowski gauge of $K$. Under these and some further assumptions on $K$, our procedures achieve the minimax rate up to small factors, depending poly-logarithmically on the dimension, while remaining computationally efficient.
We further extend our methodology to the linear regression and robust heavy-tailed settings, establishing polynomial-time near-optimal estimators when the constraint set satisfies the regularity conditions above. To the best of our knowledge, these results provide the first general framework for attaining statistically near-optimal performance under such broad geometric constraints while preserving computational tractability.
1 Introduction
Many statistical estimation problems are naturally formulated as constrained mean estimation or constrained regression problems. Given observations corrupted by stochastic noise, one seeks an estimator whose target signal is restricted to lie in a known constraint set $K\subset\mathbb{R}^{n}$. Such shape, geometric, or structural constraints arise ubiquitously in nonparametric statistics, signal processing, and machine learning—including monotonicity, convexity, sparsity, total variation, and various Lipschitz or Hölder smoothness classes (e.g., Robertson (1988); Groeneboom and Jongbloed (2014); Guntuboyina and Sen (2018); Tsybakov (2009)).
From a decision-theoretic perspective, a fundamental goal is to characterize the minimax risk over the constraint class $K$. Recent work has made significant progress on this front. In particular, Neykov (2022) derived sharp minimax rates for constrained mean estimation in the Gaussian sequence model, showing that these rates hold for essentially any convex constraint, and depend on the geometric structure of $K$. Furthermore Prasadan and Neykov (2024), relaxed the convexity assumption to a star-shaped assumption which encompasses even more constraints. In the regression setting, Prasadan and Neykov (2025a) obtained the corresponding minimax rates under analogous assumptions.
Despite this progress, a major obstacle remains: none of the procedures in these works are computationally efficient. Their constructions rely on carefully designed local packing sets adapted to the geometry of $K$. Such packings may contain exponentially many points and are typically difficult to construct or even approximate. Consequently, although these estimators achieve the optimal statistical rate, they are not implementable in polynomial-time.
Our contribution
In this paper, we develop a new framework that yields polynomial-time, near-minimax estimators for constrained mean estimation, robust mean estimation with heavy-tailed noise, and constrained linear regression. Our approach bypasses the packing-set bottleneck entirely. Instead, we build on a classical inequality of B. Carl, which relates the entropy numbers of a convex body to its Kolmogorov $n$-widths. This geometric connection allows us to translate statistical rate calculations into width-approximation problems, avoiding the need to construct packing sets.
The central technical ingredient of our work is a new algorithm for efficiently approximating optimal projections of the Kolmogorov widths of the constraint set. These approximations are sufficiently accurate to be used within an iterative estimation procedure whose risk matches the minimax rate up to small factors (which could be poly-logarithmic in the dimension e.g.).
There is, however, a price to pay: our methods require nontrivial geometric regularity of the constraint set. First, Carl’s inequality yields meaningful constants only for type-2 (see (2.1) and the discussion surrounding it for the definition of a type-2 set), origin-symmetric convex bodies. [^1] In fact, Carl’s inequality holds for all origin-symmetric convex bodies, but the resulting constants are too coarse unless the body is type-2. Second, the efficient approximation of the optimal projections of the Kolmogorov widths requires the constraint set to be type-2222Here and throughout the paper, when we say type-2 convex bodies we mean convex bodies in $\mathbb{R}^{n}$ with sufficiently small type-2 constants (see (2.1))., well-balanced, and to satisfy several additional structural conditions. While these assumptions cover a wide range of constraint classes used in modern statistics, they do exclude convex classes that are meaningful and are statistically minimax-tractable (at least in theory and some in practice). Examples of sets over which our algorithm cannot work well include rigid bodies such as the $\ell_{p}$ balls for $1\leq p<2$.
1.1 Related Work
We now review several related works. In the Gaussian sequence model setting, when the constraint set is an ellipsoid a classical result due to Pinsker Pinsker (1980) shows that linear estimators are optimal. One can implement Pinsker’s estimator in polynomial-time. A similar result holds for linear regression over ellipsoidal constraints Goldenshluger and Tsybakov (2001, 2003). Since ellipsoids are type-2 sets, our framework provides alternative polynomial-time near minimax rate optimal estimators. Notably, there exist ellipsoids over which the least squares estimator is grossly sub-optimal (see Zhang (2013) e.g.). This fact highlights the need to develop computationally efficient algorithms beyond least squares.
Zhang (2013) proposes a near minimax optimal estimator for a linear regression problem with an $\ell_{q}$ constraints on the signal for $q\leq 1$. The author’s algorithm is implementable only when $q=1$. Our work is significantly different, as we require the constraint set to be of type-2 (which the $\ell_{1}$ ball is not). Despite this major difference, there are also some similarities: both our work and Zhang (2013) use approximations of the Kolmogorov widths of the constraint set, as well as Euclidean projections. Javanmard and Zhang (2012) argue that for bounded symmetric convex polytopes, $K=\{x\in\mathbb{R}^{n}:\|Ax\|_{\infty}\leq 1\}$, defined by some matrix $A\in\mathbb{R}^{m\times n}$, the truncated series estimator [^3] A truncated series estimator uses the optimal Euclidean projection from a carefully selected Kolmogorov width, to project the data on a lower dimensional subspace. is nearly optimal for the Gaussian sequence model problem, provided that $m$ is not too large. However, the authors mention that they do not know how to efficiently implement their estimator. They also formulate a class of constraints of the type $\{x\in\mathbb{R}^{n}:\|Ax\|_{p}\leq 1\}$ for $p>0$, although they only show the near optimality result for the truncated series when $p=\infty$. Javanmard and Zhang (2012) do conjecture however (see Conjecture 7.1 therein), that the truncated series estimator remains near optimal (within $\log m$ of optimal) for sets of the type $\{x\in\mathbb{R}^{n}:\|Ax\|_{p}\leq 1\}$ for $p\geq 2$. While we cannot verify their conjecture, we do show that our estimator provably works near optimally for any well-balanced constraint of the type $K=\{x\in\mathbb{R}^{n}:\|Ax\|_{p}\leq 1\}$ for any $p\geq 2$ (we lose logarithmic factors in $m$).
Finally, we note a recent related work by Kur and Putterman (2022), which develops computationally efficient estimators for multidimensional convex and Lipschitz nonparametric regression. The estimators proposed there are fundamentally different from those studied in this paper, as they are specifically tailored to the structure of convex and Lipschitz constraints, whereas our approach is designed to address constrained estimation problems in a considerably more general setting.
1.2 Definitions and Notation
We now outline some definitions and commonly used notation.
We use a lot of absolute constants which we denote with $c,C,L,T$ and variations, whose values can change from line to line. We use $\rho_{K}$ to denote the Minkowksi gauge (also called Minkowski functional) of a convex body $K$. The gauge $\rho_{K}$ is a norm, whenever $K$ is bounded, origin-symmetric (which we will assume throughout the paper) and absorbing. We use $\|\cdot\|_{p}$ to denote the standard $\ell_{p}$ norm for $p\geq 1$. We use $\|\cdot\|$ to denote an arbitrary norm on $\mathbb{R}^{n}$. We use $\|\cdot\|_{\operatorname{op}}$, $\|\cdot\|_{F}$ and $\|\cdot\|_{*}$ to denote the operator, Frobenius and nuclear norms of a matrix respectively, and we use $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ to denote the minimum and maximum eigenvalues for symmetric positive semi-definite (psd) matrix $A$. The trace of a matrix is denoted with $\operatorname{tr}$. For two matrices $A,B\in\mathbb{R}^{m\times n}$, we put $\langle A,B\rangle=\operatorname{tr}(A^{\top}B)$ for the standard Frobenius (Hilbert-Schmidt) dot product. We write $\leq$ and $\geq$ for symmetric psd matrices to mean inequalities in the psd sense. We typically use the letter $\Pi$ to denote projections – often when indexed by a closed convex body $\Pi_{K}$ means (weak) Euclidean projection on $K$. We write $\operatorname{diag}(A)$ for a square matrix $A$ to extract its diagonal values, while $\operatorname{diag}(v)$ for a vector $v$ means a matrix with diagonal entries equal to $v$ and 0 entries otherwise. We denote the $n$-dimensional identity matrix with $I_{n}$, occasionally omitting the index when the dimension is clear from the context. We use $S^{n-1}$ for the unit Euclidean sphere in $\mathbb{R}^{n}$, while $B_{2}(x,r)$ means the closed Euclidean ball $\{y\in\mathbb{R}^{n}:\|x-y\|_{2}\leq r\}$ centered at $x$, with the shorthand $B_{2}=B_{2}(0,1)$ for the unit Euclidean ball. We will write $a\wedge b$ and $a\vee b$ to denote the minimum and maximum of two real numbers $a,b\in\mathbb{R}$. Occasionally we will use $O(1)$ and $\Omega(1)$ to denote that a quantity is upper or lower bounded by an absolute constant, respectively. The inequalities $\lesssim$ an $\gtrsim$ omit absolute constants. We write $\asymp$ if both $\lesssim$ and $\gtrsim$ hold.
We continue with the definition of Kolmogorov width which is central to this work.
Definition 1.1 (Kolmogorov width).
Let $K\subset\mathbb{R}^{n}$ be a compact set. The $k$-th Kolmogorov width of $K$ with respect to the Euclidean norm is defined as
$d_{k}(K)=\inf_{P\in\mathcal{P}_{k}}\sup_{\theta\in K}\left\lVert\theta-P\theta\right\rVert_{2},$ (1.1)
where $\mathcal{P}_{k}$ is the set of all projection operators that project a vector onto a $k$ dimensional subspace of $\mathbb{R}^{n}$.
Next we give a definition of the entropy numbers of a convex body.
Definition 1.2 (Entropy numbers).
Let $K\subset\mathbb{R}^{n}$ be a compact set. The $k$-th entropy number of $K$ with respect to the Euclidean norm is defined as
$e_{k}(K)=\inf\bigl\{\epsilon>0:\exists x_{1},\dots,x_{k}\in\mathbb{R}^{n}\text{ such that }K\subset\bigcup_{i=1}^{k}\bigl(x_{i}+\epsilon B_{2}\bigr)\bigr\}.$ (1.2)
The local entropy of a body is defined as follows.
Definition 1.3 (Local Metric Entropy).
Given a set $K$, define the $\eta$-packing number of $K$ as the maximum cardinality $M=M(\eta,K)$ of a set $\{\nu_{1},\dots,\nu_{M}\}\subset K$ such that $\|\nu_{i}-\nu_{j}\|>\eta$ for all $i\neq j$. Given a fixed sufficiently large constant $c>0$, define the local metric entropy of $K$ as $M_{K}^{\operatorname{loc}}(\eta)=\sup_{\nu\in K}M(\eta/c,B(\nu,\eta)\cap K)$. When clear from the context we drop the index $K$ from $M_{K}^{\operatorname{loc}}(\eta)$.
The local entropy plays a fundamental role in what follows as Neykov (2022); Prasadan and Neykov (2024) show it determines the minimax rate for the constraint mean estimation problem.
1.3 Organization
The paper is structured as follows. In Section 2 we give background results and we state and prove our main result conditionally on the existence of a quadratic form maximization (QFM) oracle for the convex body $K$. In Section 3 we argue that indeed such QFM oracles exist for a variety of type-2 convex bodies. In Section 4 we extend the results of Section 2 to the problem of robust heavy-tailed mean estimation. In Section 5 we discuss the extension of our main result to a linear regression setting. In Section 6 we make some remarks about sets on which our algorithm can be implemented. We close out the paper with a brief discussion of selected open problems in Section 7.
2 Background and Conditional Algorithm
Let us observe a single data point $Y=\mu+\xi$ where $\mu\in K\subset\mathbb{R}^{n}$ for some known compact convex set $K$, and $\xi=(\xi_{1},\ldots,\xi_{n})$ is a mean zero vector sub-Gaussian vector, comprised of i.i.d. components $\xi_{i}$ for $i\in[n]$. We assume that the sub-Gaussian parameter of $\xi_{i}$, $\sigma^{2}$ (or an accurate upper bound) is known to the statistician. Here by sub-Gaussian parameter we understand the smallest constant $\sigma^{2}\in\mathbb{R}_{+}$ such that for all $\lambda\in\mathbb{R}$ we have $\mathbb{E}\exp(\lambda\cdot\xi_{i})\leq\exp(\lambda^{2}\cdot\sigma^{2}/2)$. We remark that this assumption implies that $\operatorname{Var}(\xi_{i})\leq\sigma^{2}$. Denote the variance $\operatorname{Var}(\xi_{i})=\underline{\sigma}^{2}$, noting that $\underline{\sigma}^{2}\leq\sigma^{2}$.
When $\xi\sim N(0,\sigma^{2}I_{n})$, which is part of our model class, Neykov (2022) and later Prasadan and Neykov (2024) showed that when $K$ is a star-shaped set the minimax rate is given by $(\eta^{\star 2}\wedge d^{2})$ [^4] In fact, it is easy to see that $\eta^{\star}\lesssim d$ and hence the minimax rate can be written as simply $\eta^{\star 2}$., where $d=\mathrm{diam}(K)$ and $\eta^{\star}:=\eta^{\star}(K)=\sup\{\eta\geq 0:\eta^{2}/\sigma^{2}\leq\log M^{\operatorname{loc}}_{K}(\eta)\}$ with $M^{\operatorname{loc}}_{K}(\eta)$ being the local entropy of the set $K$. The goal of the present paper is to construct a polynomial-time algorithm which nearly achieves the minimax rate over a range of convex bodies $K$. Observe that we define the minimax rate with $\sigma$ rather than $\underline{\sigma}$ (in the Gaussian case which gives the lower bound, we have $\underline{\sigma}\equiv\sigma$). We will later see (see Section 4) that this can be relaxed to using $\underline{\sigma}$ in the definition of $\eta^{\star}$, but it requires having more than a single sample.
Throughout the paper, we will impose numerous requirements on the set $K$. Our first requirement is as follows. Let $K$ be symmetric about 0, so that its Minkowsi gauge $\rho_{K}(x)=\inf\{t>0:x\in tK\}$ defines a norm on $\mathbb{R}^{n}$. Our second assumption is that the type-2 constant of $K$ is “small”. Before we quantify the term “small” we define what is a type-2 constant. Let $\{\epsilon_{i}\}_{i\in[m]}$ for $m\in\mathbb{N}$ be i.i.d. Rademacher random variables. We say that $K$ has a type-2 constant $T_{2}(K)$ if for any $m\in\mathbb{N}$ and any $x_{i}\in\mathbb{R}^{n}$ we have
$\mathbb{E}_{\epsilon}\rho_{K}^{2}\bigg(\sum_{i\in[m]}\epsilon_{i}x_{i}\bigg)\leq T^{2}_{2}(K)\sum_{i\in[m]}\rho^{2}_{K}(x_{i}).$ (2.1)
In other words, $(\mathbb{R}^{n},\rho_{K})$ is a Banach space of type-2. Note that from the above definition it is clear that $T_{2}(K)\geq 1$ always. We will use this fact in several places. Going back to the “small” type-2 constant — here by small we mean that $T_{2}(X)$ is a slowly growing function of the parameters ($n,d,\sigma$). For instance $T_{2}(K)=\sqrt{\log n}$ is a small constant (in contrast to $T_{2}(K)=\sqrt{n}$ which is a large constant). Our third assumption is that we have an oracle for computing $\rho_{K}$ (which implies we have a (weak) membership oracle for $K$; see Grötschel et al. (2012) for a definition and properties of (weak) membership oracles). Our fourth assumption is that there are known constants $r,R\in\mathbb{R}_{+}$ such that $rB_{2}\subseteq K\subseteq RB_{2}$. We note here that the complexity of our algorithms depends polynomially on $\log 1/r$ and $\log R$, so in principle $r,R$ can be exponentially small/large with $n$, and hence we do not view this as a severe restriction. We will refer to such a set $K$ (i.e., a set sandwiched between two Euclidean balls) as well-balanced. Also, observe that $R=d/2$ always works, provided that $d$ is known, since the set $K$ is origin symmetric. Otherwise we must have $R\geq d/2$.
Since the space $(\mathbb{R}^{n},\|\cdot\|_{2})$ is self-dual with $T_{2}(B_{2})$ constant equal to $1$, by a classical result of B. Carl, (Carl, 1985, see Theorem 5), we have up to an absolute constant factor:
$(\prod_{k=1}^{m}d_{k}(K))^{1/m}\lesssim T_{2}(K)\inf_{k\geq 1}k^{1/m}e_{k}(K).$ (2.2)
Here $d_{k}(K)$ denotes the $k$-th Kolmogorov width of $K$ while $e_{k}(K)$ is the $k$-th entropy number (see Definitions 1.1 and 1.2). Using this inequality, we will now establish a result allowing us to find a sufficiently small Kolmogorov width.
Lemma 2.1.
Let $K$ be a type-2 convex body with type-2 constant $T_{2}(K)$, such that $rB_{2}\subseteq K\subseteq RB_{2}$ for some known $r,R$. Let $\kappa:=\kappa(K)$ be a given scalar which can depend on $K$. Let $\eta^{\star}:=\eta^{\star}(K)$ be the minimax rate defined through the equation $\eta^{\star}=\sup_{\eta}\{\eta\geq 0:\eta^{2}/\sigma^{2}\leq\log M_{K}^{\operatorname{loc}}(\eta)\}$. Assume that $\kappa^{-1/2}\tilde{d}/T_{2}(K)\geq 2\eta^{\star}$ and $\lceil\lceil\log_{c}(\kappa^{1/2}\cdot T_{2}(K))\rceil\cdot 4\eta^{\star 2}/\sigma^{2}\rceil\leq\tilde{d}^{2}/(C^{2}\sigma^{2})$ for $\tilde{d}=2R\geq d=\mathrm{diam}(K)$ and where $C>0$ is some absolute constant. Then for $m=\lceil\tilde{d}^{2}/(C^{2}\sigma^{2})\rceil\wedge n$, we have
$d_{m}(K)\lesssim\kappa^{-1/2}\cdot\tilde{d}/c,$ (2.3)
with $c$ being the constant from the definition of local entropy, which is sufficiently large.
Proof of Lemma 2.1.
First consider the case when $K$ is a singleton, i.e., $K=\{0\}$ since we are assuming it is origin-symmetric. Then the statement is clearly true since all $d_{m}(K)\equiv 0$ for all $m\in[n]$. Furthermore if $\sigma=0$, $m=n$ and thus (2.3) continues to hold. Next, consider the non-trivial case when $K$ contains at least two points and $\sigma>0$. It is easy to see that the former implies that for small $\delta$, $\log M^{\operatorname{loc}}(\delta)\geq\log 2c>1$. This in turn implies, $\eta^{\star}>0$.
Observe now that for any $\delta\geq 2\eta^{*}$ we have $\log M^{\operatorname{loc}}(\delta)\leq\log M^{\operatorname{loc}}(2\eta^{\star})\leq 4\eta^{\star 2}/\sigma^{2}$ (see Neykov (2022) Lemma II.8 where it is shown that for convex sets the map $\delta\mapsto\log M^{\operatorname{loc}}(\delta)$ is non-increasing). Set $\kappa^{-1/2}\cdot\tilde{d}/T_{2}(K)=\Upsilon$. Under our assumptions we have that $\Upsilon\geq 2\eta^{*}$. Next, by Lemma 3 of Yang and Barron (1999) we know that
$\log M(\Upsilon c^{j-1})-\log M(\Upsilon c^{j})\leq\log M^{\operatorname{loc}}(\Upsilon c^{j})\leq 4\eta^{\star 2}/\sigma^{2},$ (2.4)
for $j=0,1,\ldots$. Here recall from the definition of local entropy that we use $M(\delta)$ to denote the $\delta$ packing of $K$. Note that for $j\geq\lceil\log_{c}(\kappa^{1/2}\cdot T_{2}(K))\rceil$ we have $\log M(\Upsilon c^{j})=0$. Thus upon summing all the inequalities for $j=0,\ldots,\lceil\log_{c}(\kappa^{1/2}\cdot T_{2}(K))\rceil$ we have
$\log M(\Upsilon c^{-1})\leq\lceil\log_{c}(\kappa^{1/2}\cdot T_{2}(K))\rceil 4\eta^{\star 2}/\sigma^{2},$
This implies that for $k=M(\Upsilon c^{-1})\leq\exp(\lceil\log_{c}(\kappa^{1/2}\cdot T_{2}(K))\rceil 4\eta^{\star 2}/\sigma^{2})$, we have $e_{k}(K)\leq\kappa^{-1/2}\cdot\tilde{d}/(T_{2}(K)c)$. Hence by (2.2) and since the Kolmogorov widths are decreasing in $m$, we have
$d_{m}(K)\leq(\prod_{k=1}^{m}d_{k}(K))^{1/m}\lesssim T_{2}(K)\exp(\lceil\log_{c}(\kappa^{1/2}\cdot T_{2}(K))\rceil 4\eta^{\star 2}/\sigma^{2}/m)\kappa^{-1/2}\cdot\tilde{d}/(T_{2}(K)c).$ (2.5)
Next select $m=\lceil\tilde{d}^{2}/(C^{2}\sigma^{2})\rceil\wedge n$, where $\tilde{d}=2R$ is an upper bound on $d$. Since both $R$ and $\sigma$ are assumed to be known, this is a legitimate choice of $m$. By assumption $\lceil\lceil\log_{c}(\kappa^{1/2}\cdot T_{2}(K))\rceil\cdot 4\eta^{\star 2}/\sigma^{2}\rceil\leq\tilde{d}^{2}/(C^{2}\sigma^{2})$, it is clear by (2.5) that:
$d_{m}(K)\lesssim\kappa^{-1/2}\cdot\tilde{d}/c,$
where we note that in case $m=n$ the above inequality is trivial as $d_{n}(K)=0$. ∎
Unfortunately, we do not know the optimal projection that yields a small Kolmogorov width. In addition, even calculating the Kolmogorov width $d_{m}(K)$ precisely in general is likely NP hard – see, for instance, Brieden (2002) where the author shows that calculating $d_{0}(K)$ for (even symmetric) convex polytopes cannot be computed in polynomial-time within a factor of $1.090$, unless P=NP. Thus, the fact (2.3) appears useless at first glance. However, since we are interested in having a minimax optimal estimator up to “small” factors, we do not need to find the precise optimal Kolmogorov projection. Taking inspiration from Varadarajan et al. (2007), we propose to solve the following semi-definite program (SDP) relaxation of the Kolmogorov width problem
$\mathop{\mathrm{argmin}}_{X}\max_{p\in K}p^{\top}Xp,$ (2.6)
$\mbox{subject to }\operatorname{tr}(X)=n-m,$
$~~~~~~~~~~~~~~~0\leq X\leq I,X^{\top}=X$
As observed in Varadarajan et al. (2007), the optimal value $X^{\star}$ that this SDP returns, satisfies $p^{\top}X^{\star}p\leq d^{2}_{m}(K)$ for all $p\in K$. The proof of this fact is elementary, and we include a proof below for convenience of the reader. The proof also helps motivate program (2.6).
Lemma 2.2.
The solution to the above SDP problem, $X^{\star}$, satisfies: $p^{\top}X^{\star}p\leq d^{2}_{m}(K)$ for all $p\in K$.
Proof of Lemma 2.2.
Let $P_{n-m}=I-P_{m}$ be the optimal Kolmogorov with projection, i.e.,
$\sup_{p\in K}\|(I-P_{m})p\|_{2}^{2}=d^{2}_{m}(K).$
This is equivalent to $\sup_{p\in K}p^{\top}P_{n-m}p\leq d^{2}_{m}(K)$. Note that $P_{n-m}$ is a feasible point to the SDP hence the result follows. ∎
Notably, solving the SDP requires maximizing a quadratic form over a convex body (cf. (2.6)), a formidable task on its own. The authors of Varadarajan et al. (2007) do not encounter this issue, since their body $K$ is a convex polytope with finitely many vertices, making the quadratic maximization tractable (the maximum is attained at a vertex and can be found by enumerating all vertices). In our setting, no such guarantee holds. We will now assume that there exists an oracle $\mathcal{O}_{K}$ which takes a symmetric psd matrix $X$ and outputs a vector $\mathcal{O}_{K}(X)\in K$ such that
$\mathcal{O}_{K}(X)^{\top}X\mathcal{O}_{K}(X)\geq\kappa^{-1}\max_{p\in K}p^{\top}Xp,$ (2.7)
for some “small” [^5] Here by small we mean that $\kappa(K)$ is a constant slowly growing with $n$. constant $\kappa=\kappa(K)\geq 1$ with probability at least 1q1-q for some $q\in\mathbb{R}$ with $\log q^{-1}$ being a parameter involved in the number of iterations needed to compute the oracle, whose value will be specified later (see Section 3). In other words, we suppose there exists an approximate solution to the quadratic maximization problem. Later we will see that there is a rich class of type-2 bodies for which such an oracle indeed exits. In order to state our algorithm for approximately solving the SDP, we consider the projection of a vector $v\in\mathbb{R}^{n}$ onto the convex set
$S_{k}=\{\,w\in\mathbb{R}^{n}:0\leq w_{i}\leq 1,\ \sum_{i=1}^{n}w_{i}=k\,\},$
where $k\in\{0,\ldots,n\}$ is given. The projection is defined as
$\Pi_{S_{k}}(v)=\arg\min_{w\in S_{k}}\frac{1}{2}\|w-v\|_{2}^{2}.$
$\Pi_{S_{k}}(v)$ can be implemented efficiently (see Algorithm 2). Given the oracle $\mathcal{O}_{K}$ and the projection $\Pi_{S_{k}}(\cdot)$, we may approximately solve the SDP with projected sub-gradient descent. We formalize the algorithm below:
[…]
1 SDP with projected sub-gradient descent
0: Convex body $K$, value $m\in\{0,\ldots,n\}$, oracle $\mathcal{O}_{K}$ for solving quadratic maximization over $K$, a tuning parameter $C\in\mathbb{R}_{+}$, $M=4(n-m)C^{4}\kappa^{2}$ maximal number of iterations, $\gamma=1/(\kappa C^{2})\in\mathbb{R}$ is the learning rate.
0: A symmetric psd matrix $X^{\star\star}$
1: Set $X=\operatorname{diag}(((n-m)/n,(n-m)/n,\ldots,(n-m)/n))$
2: best = $\infty$
3:for j = 1, …, M do
4:  if best $\mathcal{O}_{K}(X)^{\top}X\mathcal{O}_{K}(X)$ then
5:   $X^{\text{best}}=X$
6:   best = $\mathcal{O}_{K}(X)^{\top}X\mathcal{O}_{K}(X)$
7:  end if
8:  Set $\gamma_{j}=\gamma/\|\mathcal{O}_{K}(X)\|_{2}^{2}$
9:  $X=X-\gamma_{j}\mathcal{O}_{K}(X)\mathcal{O}_{K}(X)^{\top}$
10:  Eigendecompose $X=U\Lambda U^{\top}$
11:  $v=\Pi_{S_{n-m}}(\operatorname{diag}(\Lambda))$
12:  $X=U\operatorname{diag}(v)U^{\top}$
13:end for
14:return $X^{\star\star}:=X^{\text{best}}$
Next we write down the sub-routine for $\Pi_{S_{n-m}}$ used in Algorithm 2 as Algorithm 2. We give the justification of Algorithms 2 and 2 below.
[…]
2 Projection onto the capped simplex $S_{k}=\{w:0\leq w_{i}\leq 1,\sum_{i}w_{i}=k\}$
0: Vector $v\in\mathbb{R}^{n}$, integer $k\in\{0,\ldots,n\}$
0: Projected vector $w^{\star}=\arg\min_{w\in S_{k}}\|w-v\|_{2}^{2}$
1: Define breakpoints $\mathcal{B}=\{v_{i}-1,v_{i}:i=1,\ldots,n\}\cup\{+\infty\}$
2: Sort the breakpoints in increasing order
3:for each interval $b_{j},b_{j+1})$ between consecutive breakpoints do
4:  Let $I_{0}=\{i:v_{i}-b_{j}\geq 1\}$, $I_{1}=\{i:0<v_{i}-b_{j}<1\}$
5:  if $|I_{1}|==0$ then
6:   if $|I_{0}|==k$ then
7:    Set $\theta=b_{j}$
8:    break
9:   end if
10:   next
11:  end if
12:  Compute
$\theta=\frac{\sum_{i\in I_{1}}v_{i}+|I_{0}|-k}{|I_{1}|}$
13:  if $\theta\in[b_{j},b_{j+1})$ then
14:   break
15:  end if
16:end for
17: Return $w_{i}^{\star}=\min\{1,\max\{0,v_{i}-\theta\}\}$ for all $i$
We start by justifying Algorithm 2 which finds $\Pi_{S_{k}}v$. Specifically, we have:
Lemma 2.3 (Projection onto the capped simplex).
Let $v\in\mathbb{R}^{n}$ and $k\in\{0,\ldots,n\}$. Then the unique minimizer of
$\min_{w\in S_{k}}\frac{1}{2}\|w-v\|_{2}^{2}$
is given componentwise by
$w_{i}^{\star}=\min\{1,\max\{0,v_{i}-\theta^{\star}\}\},$
where the scalar $\theta^{\star}$ is the unique solution of
$\sum_{i=1}^{n}\min\{1,\max\{0,v_{i}-\theta^{\star}\}\}=k.$
In addition, observe that $w_{i}$ are ordered in the same way, in other words $(v_{i}-v_{j})(w_{i}-w_{j})\geq 0$ for any $i,j$.
The proof of Lemma 2.3 is deferred to the appendix. It is clear that this justifies Algorithm 2. We now argue that the projection of a matrix, on the set of symmetric psd matrices which are smaller than the identity matrix in the psd sense and have fixed trace, is given by a simple spectral decomposition plus a projection on the capped simplex.
Lemma 2.4.
Let $X$ be a symmetric psd matrix. Then the projection of $X$ onto the set $\{W\text{ symmetric psd}:0\leq W\leq I,\operatorname{tr}(W)=k\}$ in the Frobenius norm (i.e., $W^{\star}=\mathop{\mathrm{argmin}}_{W}\|X-W\|_{F}^{2}$) is given by
$W^{\star}=U\operatorname{diag}(\Pi_{S_{k}}(\operatorname{diag}(\Lambda)))U^{\top},\text{ where }X=U\Lambda U^{\top},$
is the spectral decomposition of $X$.
Proof.
Let $W^{\star}$ be the minimizer of $\|X-W^{\star}\|_{F}^{2}$ for W{W $^{\star}\in\{W\text{ symmetric psd}:0\leq W\leq I,\operatorname{tr}(W)=k\}$. Suppose $W^{\star}=U^{\star}\Lambda^{\star}U^{\star\top}$, where $\Lambda^{\star}$ is a diagonal matrix sorted in descending way (same as sorting as the diagonal entries of $\Lambda$). Let $\lambda=\operatorname{diag}(\Lambda)$ and $\lambda^{\star}=\operatorname{diag}(\Lambda^{\star})$. We have
$\|X-W^{\star}\|_{F}^{2}=\|X\|_{F}^{2}-2\operatorname{tr}(X^{\top}W)+\|W\|_{F}^{2}=\|\lambda\|_{2}^{2}-2\operatorname{tr}(X^{\top}W)+\|\lambda^{\star}\|_{2}^{2}.$
Note that
$\operatorname{tr}(X^{\top}W)=\sum_{i}\sum_{j}(u_{i}^{\top}u_{j}^{\star})^{2}\lambda_{i}\lambda^{\star}_{j}=\lambda^{\top}A\lambda^{\star},$
where $A=\{A_{ij}\}_{i,j\in[n]}=(u_{i}^{\top}u_{j}^{\star})^{2}$ is a doubly stochastic matrix (by the orthonormality of $U$ and $U^{\star}$). By Birkhoff–von Neumann’s theorem $A=\sum_{i\in[k]}\alpha_{i}P_{i}$ for some permutation matrices $P_{i}$, $k\in\mathbb{N}$ and $\sum_{i\in[k]}\alpha_{i}=1$. Thus
$\operatorname{tr}(X^{\top}W)=\sum_{i\in[k]}\alpha_{i}\lambda P_{i}\lambda^{\star}\leq\lambda^{\top}\lambda^{\star},$
where the last inequality comes from the rearrangement inequality since both $\lambda$ and $\lambda^{\star}$ are vectors sorted in descending order.
We conclude that
$\|X-W^{\star}\|_{F}^{2}\geq\|\lambda\|_{2}^{2}-2\lambda^{\top}\lambda^{\star}+\|\lambda^{\star}\|_{2}^{2}=\|\lambda-\lambda^{\star}\|_{2}^{2}.$
The above inequality is also clearly an equality whenever $U^{\star}=U$. Since the projection of $\lambda$ onto the set $S_{k}$ is $\Pi_{S_{k}}(\lambda)$ as we showed in Lemma 2.3 the result follows.
The following theorem establishes the validity of Algorithm 2.
Theorem 2.5.
Given as input a convex body $K$, a value $m$, and an oracle $\mathcal{O}_{K}$ satisfying (2.7) with probability $q=\frac{M^{-1}}{1+\tilde{d}^{3}/\sigma^{3}}$, Algorithm 2 outputs a matrix $X^{\star\star}$ which up to an absolute constant factors satisfies
$\sup_{p\in K}p^{\top}X^{\star\star}p\lesssim\tilde{d}^{2}/c^{2},$
with probability at least $1-(1+\tilde{d}^{3}/\sigma^{3})^{-1}$.
Remark 2.6.
Observe that since $M=4(n-m)C^{4}\kappa^{2}$ and the oracle works with polynomially many iterations in $\log q^{-1}$ (see Section 3), if $\kappa$ is a “small” constant (i.e., $\kappa\ll n$), and in addition $\log\tilde{d}/\sigma$ is polynomial in $n$, we are guaranteed to have $\log[M(1+\tilde{d}^{3}/\sigma^{3})]$ is polynomial in $n$, so that our oracle algorithm will run in polynomial time.
Proof of Theorem 2.5.
We first establish the high probability bound. Since we have $M$ iterations, and the oracle succeeds with probability $1-M^{-1}/(1+\tilde{d}^{3}/\sigma^{3})$ every time, by the union bound it will fail with probability smaller than $1/(1+\tilde{d}^{3}/\sigma^{3})$, which establishes the claim. Let us now show the guarantee assuming that the oracle never fails.
Let us begin by observing that the function $f(X):=\sup_{p\in K}p^{\top}Xp=\sup_{p\in K}\langle X,pp^{\top}\rangle$ is convex as a maximum of linear functions. It is also well known that $p^{\star}p^{\star\top}$ is a sub-gradient of this function, where $p^{\star}=\mathop{\mathrm{argmax}}_{p\in K}\langle X,pp^{\top}\rangle$. Next, note that by our definition of the oracle, the vector $\mathcal{O}_{K}(X)$ is near sub-gradient of the map in the following sense:
$f(X)=\sup_{p\in K}\langle X,pp^{\top}\rangle\leq\langle X,\mathcal{O}_{K}(X)\mathcal{O}_{K}(X)^{\top}\rangle+(1-\kappa^{-1})\cdot f(X)$ (2.8)
Let $X^{(k)}$ be the matrix $X$ in the $k$-th iteration, and put $g^{(k)}=\mathcal{O}_{K}(X^{(k)})\mathcal{O}_{K}(X^{(k)})^{\top}$. Recall that $X^{\star}$ is the true optimum, which satisfies by Lemma 2.2 that $\sup_{p\in K}p^{\top}X^{\star}p\leq d^{2}_{m}(K)$ for all $p\in K$. Observe that we have
$f(X^{(k)})-f(X^{\star})\leq\langle g^{(k)},X^{(k)}-X^{\star}\rangle+(1-\kappa^{-1})\cdot f(X^{(k)}),$
where we used (2.8) and the fact that $\langle g^{(k)},X^{\star}\rangle\leq f(X^{\star})$. We will now repeat the original proof of the sub-gradient descent due to Shor Shor (2012); Boyd and Vandenberghe (2004), using the near sub-gradient property above. Since Euclidean projection on a convex set shrinks the distance to any point in the set, we have
$\|X^{(k+1)}-X^{\star}\|_{F}^{2}$ $\leq\|X^{(k)}-X^{\star}\|_{F}^{2}+2\gamma_{k}\langle g^{(k)},X^{\star}-X^{(k)}\rangle+\gamma_{k}^{2}\|g^{(k)}\|_{F}^{2}$
$\leq\|X^{(k)}-X^{\star}\|_{F}^{2}+2\gamma_{k}(f(X^{\star})-\kappa^{-1}\cdot f(X^{(k)}))+\gamma_{k}^{2}\|g^{(k)}\|_{F}^{2}.$
Thus after $M$ iterations we have ,
$2\sum_{k=1}^{M}\gamma_{k}\min_{k\in[M]}(\kappa^{-1}\cdot f(X^{(k)})-f(X^{\star}))$ $\leq 2\sum_{k=1}^{M}\gamma_{k}(\kappa^{-1}\cdot f(X^{(k)})-f(X^{\star}))$
$\leq\|X^{(1)}-X^{\star}\|_{F}^{2}+\sum_{k=1}^{M}\gamma_{k}^{2}\|g^{(k)}\|_{F}^{2}.$
Or after dividing by $\sum_{k=1}^{M}\gamma_{k}$ we obtain:
$\min_{k\in[M]}(\kappa^{-1}\cdot f(X^{(k)})-f(X^{\star}))\leq\frac{\|X^{(1)}-X^{\star}\|_{F}^{2}+\sum_{k=1}^{M}\gamma_{k}^{2}\|g^{(k)}\|_{F}^{2}}{2\sum_{k=1}^{M}\gamma_{k}}.$
Next we note that $\|g^{(k)}\|_{F}\leq d_{0}(K)^{2}\leq d^{2}$ for any $k$ (since $\mathcal{O}_{K}(X^{(k)})\in K$). Furthermore, $\|X^{(1)}\|_{F}^{2}=\sum_{i\in[n]}\lambda^{2}_{i}(X^{(1)})\leq\sum_{i\in[n]}\lambda_{i}(X^{(1)})\leq n-m$, and a similar logic shows that $\|X^{\star}\|_{F}^{2}\leq(n-m)$. Hence $\|X^{(1)}-X^{\star}\|_{F}^{2}\leq 4(n-m)$. Hence upon selecting, $\gamma_{k}=\gamma/\|g^{(k)}\|_{F}$ we obtain
$\min_{k\in[M]}(\kappa^{-1}\cdot f(X^{(k)})-f(X^{\star}))\leq\frac{4(n-m)+M\gamma^{2}}{2M\gamma/d^{2}}.$
Therefore taking into account that $\gamma=1/(\kappa\cdot C^{2})$, it suffices to take $M\geq 4(n-m)C^{4}\kappa^{2}$. We thus conclude that the smallest $f(X^{(\text{best})})=\min_{j\in[M]}f(X^{(k)})$ satisfies
$f(X^{(\text{best})})$ $\leq\kappa\cdot(\kappa^{-1}\cdot\tilde{d}^{2}/C^{2}+f(X^{\star}))\leq\kappa\cdot(\kappa^{-1}\cdot\tilde{d}^{2}/C^{2}+d_{m}(K)^{2})\leq\kappa\cdot(\kappa^{-1}\cdot\tilde{d}^{2}/C^{2}+L\cdot\kappa^{-1}\cdot\tilde{d}^{2}/c^{2})$
$\lesssim\tilde{d}^{2}/c^{2},$
where we used $\tilde{d}\geq d$, which is what we aimed to show.
Once we have obtained the matrix $X^{\star\star}=V\Lambda V^{\top}=\sum_{i\in[n]}\lambda_{i}v_{i}v_{i}^{\top}$, our algorithm proceeds with constructing the matrix $A^{\star\star}=(I-X^{\star\star})^{1/2}=\sum_{i\in[n]}\sqrt{1-\lambda_{i}}v_{i}v_{i}^{\top}$. Consider now the vector $A^{\star\star}Y$. It is equal to
$A^{\star\star}Y=A^{\star\star}\mu+A^{\star\star}\xi.$
In order to estimate $A^{\star\star}\mu$, we will use $A^{\star\star}Y$. We have the following lemma justifying that this is not a bad choice.
Lemma 2.7.
Supopse $\tilde{d}\gtrsim\sigma$. Then, the estimator $A^{\star\star}Y$ of $A^{\star\star}\mu$ satisfies
$\|A^{\star\star}Y-A^{\star\star}\mu\|_{2}^{2}\leq\tilde{d}^{2}/T^{2},$
with probability at least $1-\exp(-\tilde{d}^{2}/(\sigma^{2}C^{{}^{\prime}2}))-(1+\tilde{d}^{3}/\sigma^{3})^{-1}$.
Proof.
Let us first condition on the event that our oracle did not fail (which happens with probability at least $1-(1+\tilde{d}^{3}/\sigma^{3})^{-1}$).
Applying the Hanson-Wright inequality Rudelson and Vershynin (2013), we obtain:
$\mathbb{P}(\|A^{\star\star}\xi\|_{2}^{2}-\underline{\sigma}^{2}\operatorname{tr}(A^{\star\star 2})\geq t)\leq\exp(-ct^{2}/(\sigma^{4}\|A^{\star\star 2}\|^{2}_{F})\wedge t/(\sigma^{2}\|A^{\star\star 2}\|_{\operatorname{op}}))$ (2.9)
Note that $\|A^{\star\star 2}\|_{\operatorname{op}}\leq 1$. In addition $\|A^{\star\star 2}\|^{2}_{F}=\operatorname{tr}((I-X^{\star\star})^{2})\leq\operatorname{tr}(I-X^{\star\star})=m$.
Hence we can select $t=\tilde{d}^{2}/c^{2}\gtrsim m\sigma^{2}$ to claim that (2.9) holds with high probability, i.e., probability at least $1-\exp(-c^{\prime}\tilde{d}^{4}/(\sigma^{4}m)\wedge\tilde{d}^{2}/(c^{2}\sigma^{2}))$. Comparing $\tilde{d}^{4}/(\sigma^{4}m)\gtrsim\tilde{d}^{2}/\sigma^{2}$ yields $\tilde{d}^{2}\gtrsim\sigma^{2}m$ which is true as $m=\lceil\tilde{d}^{2}/(C^{2}\sigma^{2})\rceil\wedge n$ and $\tilde{d}\gtrsim\sigma$. This implies $c^{\prime}\tilde{d}^{4}/(\sigma^{4}m)\wedge\tilde{d}^{2}/(c^{2}\sigma^{2})\asymp\tilde{d}^{2}/(C^{{}^{\prime}2}\sigma^{2})$ for some $C^{\prime}>0$. So that the bound is always $1-\exp(-\tilde{d}^{2}/(C^{{}^{\prime}2}\sigma^{2}))$. Since the mean of $\|A^{\star\star}\xi\|_{2}^{2}$ is $\underline{\sigma}^{2}\operatorname{tr}(A^{\star\star 2})=\underline{\sigma}^{2}m\lesssim\sigma^{2}\tilde{d}^{2}/(C^{2}\sigma^{2})$ we complete the proof. ∎
Remark 2.8.
Lemma 2.7 is the reason why we assumed that the noise $\xi$ has independent components, so that we can apply the Hanson-Wright inequality. We note that the formula for the minimax rate remains unchanged even if the noise has dependent entries Prasadan and Neykov (2024). In Section 4 we will see that if we have more samples the noise can be arbitrary provided that its covariance matrix has a bounded operator norm.
Our main algorithm starts at 0 and works by iteratively refining the estimate arguing that at each step we have better and better approximation of $\mu$. These approximations will be shown to decay exponentially as the iteration proceeds, and in that sense the following lemma is very useful. Assuming that we have ”trapped” $\mu$ in an Euclidean ball of certain radius we demonstrate how we can find an even better estimator of $\mu$ provided that the conditions of Lemma 2.1 are satisfied.
Lemma 2.9.
Suppose we have an estimator of $\mu$ such that $\|\hat{\mu}_{j}-\mu\|_{2}\leq\tilde{d}_{j}$ happens with high probability, for some $\tilde{d}_{j}\in\mathbb{R}_{+}$ such that $\tilde{d}_{j}\geq 2r\vee C\sigma$. Let $K_{(j)}=K\cap B(0,\tilde{d}_{j}/2)$. Obtain $m$ and the matrix $A_{j}^{\star\star}$ from Lemma 2.1 and Algorithm 2 for the set $K_{(j)}$ [^6] We explain later that the oracle $\mathcal{O}_{K_{(j)}}$ can also be run in polynomial-time, provided that the oracle $\mathcal{O}_{K}$ can be run in polynomial-time., assuming that the conditions of Lemma 2.1 hold. Then the estimator $\tilde{\mu}_{j+1}=(A^{\star\star}_{j}Y-A_{j}^{\star\star}\hat{\mu}_{j})/2$ of $(\mu-\hat{\mu}_{j})/2$ used in Lemma 2.7 satisfies:
$\mathbb{P}(\|\tilde{\mu}_{j+1}-(\mu-\hat{\mu}_{j})/2\|_{2}\geq\tilde{d}_{j}/L,\|\hat{\mu}_{j}-\mu\|_{2}\leq\tilde{d}_{j})\leq\exp(-\tilde{d}_{j}^{2}/(C^{{}^{\prime}2}\sigma^{2}))+\frac{1}{1+\tilde{d}^{3}/\sigma^{3}}.$
for some absolute constant $L$.
Proof.
On the event $\|\hat{\mu}_{j}-\mu\|_{2}\leq\tilde{d}_{j}$ we have that, $\nu=(\mu-\hat{\mu}_{j})/2\in K_{(j)}=K\cap B(0,\tilde{d}_{j}/2)$. Recall that $A_{j}^{\star\star}=(I-X_{j}^{\star\star})^{1/2}$, and let $X_{j}^{\star\star}=V\Lambda V^{\top}$ be the eigendecomposition of $X_{j}^{\star\star}$.
$\|\nu-A_{j}^{\star\star}\nu\|_{2}^{2}$ $=\|\nu-\sum_{i}\sqrt{1-\lambda_{i}}v_{i}(v_{i}^{\top}\nu)\|^{2}_{2}=\sum_{i}(1-\sqrt{1-\lambda_{i}})^{2}(v_{i}^{\top}\nu)^{2}$
$=\sum_{i}\frac{\lambda_{i}^{2}}{(1+\sqrt{1-\lambda_{i}})^{2}}(v_{i}^{\top}\nu)^{2}\leq\sum_{i}\lambda_{i}(v_{i}^{\top}\nu)^{2}=\nu^{\top}X_{j}^{\star\star}\nu\lesssim(\tilde{d}_{j}/c)^{2},$
where we used $0\leq\lambda_{i}\leq 1$. Next, $\|\tilde{\mu}_{j+1}-(\mu-\hat{\mu}_{j})/2\|_{2}\leq\|A_{j}^{\star\star}Y/2-A_{j}^{\star\star}\mu/2\|_{2}+\|A_{j}^{\star\star}(\mu-\hat{\mu}_{j})/2-(\mu-\hat{\mu}_{j})/2\|_{2}\lesssim\tilde{d}_{j}/(2T)+(\tilde{d}_{j}/c)\leq\tilde{d}_{j}/L$, with probability at least $1-\exp(-\tilde{d}_{j}^{2}/(C^{{}^{\prime}2}\sigma^{2}))-(1+\tilde{d}^{3}/\sigma^{3})$ by Lemma 2.7. This completes the proof. ∎
Finally we can describe the iteration of our algorithm.
[…]
3 Near-optimal Estimation Algorithm
0: Convex body $K$, $r,R\in\mathbb{R}_{+}$ such that $rB_{2}\subseteq K\subseteq RB_{2}$, observation $Y=\mu+\xi$, constant $\tilde{L}=L/(\sqrt{3}+1)$ where $L$ is from Lemma 2.9; We assume $\tilde{L}>2(\sqrt{3}+1)$ [^7] This assumption can be met upon selecting $C$ from the definition of $m$ and $c$ from the definition of local entropy sufficiently large.
0: Estimator $\hat{\mu}$ of $\mu$
1: Set $\tilde{d}_{1}=2R$, $K_{(1)}:=K$, number of iterations $M=\lceil\log_{\tilde{L}/(2(\sqrt{3}+1))}R/r\rceil$
2: Set $\hat{\mu}_{1}:=0$
3:if $\sigma\leq r/\sqrt{n}$ then
4:  return $Y$;
5:end if
6:for j = 1, …, M do
7:  Find $m=\lceil\tilde{d}_{j}^{2}/(C^{2}\sigma^{2})\rceil\wedge n$, and calculate the matrix $A_{j}^{\star\star}$ for the set $K_{(j)}$
8:  Find $\tilde{\mu}_{j+1}=(A_{j}^{\star\star}Y-A_{j}^{\star\star}\hat{\mu}_{j})/2$
9:  Set $\bar{\mu}_{j+1}=\Pi_{K_{(j)}}\tilde{\mu}_{j+1}$ [^8] Here we set $\epsilon=\tilde{d}_{j}/L$ in Lemma A.1 when evaluating the weak projection.
10:  Lift back to $K$: $\hat{\mu}_{j+1}=\Pi_{K}(2\bar{\mu}_{j+1}+\hat{\mu}_{j})$ [^9] Here we set $\epsilon=2\tilde{d}_{j}/\tilde{L}$ in Lemma A.1 when evaluating the weak projection.
11:  Set $\tilde{d}_{j+1}=2(\sqrt{3}+1)\tilde{d}_{j}/\tilde{L}$
12:  Set $K_{(j+1)}:=K\cap B_{2}(0,\tilde{d}_{j+1}/2)$
13:  if $\tilde{d}_{j+1}\leq(2r)\vee(C\sigma)$ then
14:   return $\hat{\mu}:=\hat{\mu}_{j+1}$
15:  end if
16:end for
17:return $\hat{\mu}:=\hat{\mu}_{M+1}$
We make several comments about the algorithm. Note that, since we can evaluate the Minkowski gauge of $K$, we have a (weak) separation oracle for $K$ Grötschel et al. (2012), which in turn enables (weak) projection in polynomial-time (see Theorem 2.5.9 in Dadush (2012), see also Lee et al. (2018)). The properties of this weak projection are summarized in Lemma A.1, where it is observed that it can output a vector that is in $K$ which is closer (up to an absolute constant factor) to any vector in the set $K$ than the original vector. When we write $\Pi_{K}$ we mean using this weak projection. In addition, for any $j$ the set $K_{(j)}$ also admits weak projection. This is because the Minkowski gauge of $K_{(j)}$ satisfies
$\rho_{K_{(j)}}(x)=\rho_{K}(x)\vee(2/\tilde{d}_{j}\cdot\|x\|_{2}),$ (2.10)
and can therefore be explicitly evaluated. In addition $K_{(j)}$ is well balanced since $rB_{2}\subseteq K_{(j)}\subseteq\tilde{d}_{j}/2B_{2}$. Thus when we write $\Pi_{K_{(j)}}$ we mean the weak projection on $K_{(j)}$. Since the projections are “weak” they need not shrink distance to any point in $K$. Hence the need for scaling factors: $(\sqrt{3}+1)$ and $2(\sqrt{3}+1)$ in our algorithm.
We remark that our estimator is not proper, because when $r>\sigma\sqrt{n}$ the estimator $Y$ need not be in $K$. This can be remedied by outputting $\Pi_{K}Y$ in the case when $r>\sigma\sqrt{n}$, with $\epsilon=\sigma\sqrt{n}$ in Lemma A.1 or setting $r=r\wedge\sqrt{n}\sigma$ which circumvents this issue altogether.
Next, to ensure that the matrix $A_{j}^{\star\star}$ is calculable, we need to show the exsitance of the oracle $\mathcal{O}_{K_{(j)}}$ for any $j=1,\ldots,M$. We will do so in the following section, Section 3, but for now we will show that if the original set $K$ is type-2, the sets $K_{(j)}$ remain type-2. This is so since its Minkowski gauge satisfies (2.10), and is thus type-2 with constant at most
$T_{2}(K_{(j)})\leq\sqrt{2}\max(T_{2}(K),T_{2}(B_{2}(0,\tilde{d}_{j}/2)))=\sqrt{2}\max(T_{2}(K),1)=\sqrt{2}T_{2}(K).$
Furthermore, as $K_{(j)}\subseteq K$, it is clear that the local entropy $\log M^{\operatorname{loc}}_{K_{(j)}}(\delta)\leq\log M^{\operatorname{loc}}_{K}(\delta)$ for any $\delta$. What is more, the latter implies that we have $\eta^{\star}(K_{(j)})\leq\eta^{\star}(K)$.
Before we go on to show that our algorithm is nearly minimax optimal we first show the following lemma, justifying why one can output $Y$ when $\sigma$ is small.
Lemma 2.10.
Let $\sigma\gtrsim r/\sqrt{n}$. Then the minimax rate $\eta^{\star}\gtrsim r$. On the other hand if $\sigma\lesssim r/\sqrt{n}$ we have $\eta^{\star}\asymp\sqrt{n}\sigma$.
Proof.
Recall that the minimax rate $\eta^{\star}$ satisfies $\eta^{\star}=\sup\{\eta\geq 0:\eta^{2}/\sigma^{2}\leq\log M^{\operatorname{loc}}(\eta)\}$. Observe that for $\eta\lesssim r$ we have $\log M^{\operatorname{loc}}(\eta)\asymp n$. On the other hand since $\sigma\gtrsim r/\sqrt{n}$, the ratio $(\eta/\sigma)^{2}\lesssim n$. Thus $\eta^{\star}\geq\eta$ by definition. Since $\eta$ can be taken as $\eta\asymp r$ the proof follows.
For the second part observe that since $\sigma\lesssim r/\sqrt{n}$ then $\sqrt{n}\sigma\lesssim r$ and therefore $\log M^{\operatorname{loc}}(\sqrt{n}\sigma)\asymp n$. Thus for $\eta\asymp\sqrt{n}\sigma$ it follows that $\eta^{2}/\sigma^{2}\asymp\log M^{\operatorname{loc}}(\eta)$. On the other hand, $\eta^{\star}$ cannot be larger than $\sqrt{n}\sigma$ since otherwise $\eta^{\star 2}/\sigma^{2}\gtrsim n\gtrsim\log M^{\operatorname{loc}}(\eta^{\star})$. ∎
We now state and prove the main result of the paper:
Theorem 2.11.
Suppose that $T_{2}(K)$ and $\max_{i\in[M]}\kappa(K_{(k)})$ are “small” factors, and that $\log(R/r)$ is polynomial in $n$. Then there exists a polynomial up to logarithmic factors function $\Upsilon(K)=g(T_{2}(K),\max_{i\in[M]}\kappa(K_{(k)}))$ (see the proof for details) such that the output $\hat{\mu}$ of Algorithm 2 satisfies
$\mathbb{E}\|\hat{\mu}-\mu\|_{2}^{2}\lesssim\Upsilon(K)\eta^{\star 2}(K),$
and is thus nearly minimax optimal.
Proof.
Clearly, on the first step,
$\|\mu-0\|_{2}\leq 2R=\tilde{d}_{1}.$
We will now explicitly write out the first couple of iterations of our algorithm, to ease the reader in the inductive logic. Every time, we will condition on the event $\|\tilde{\mu}_{i+1}-(\mu-\hat{\mu}_{i})/2\|_{2}\leq\tilde{d}_{i}/L$. In the end we will use the union bound using logic similar to Lemma B.5 of Prasadan and Neykov (2024) to place ourselves on all of these events.
Apply Lemma 2.1 to the set $K_{(1)}$. We know that either one of $\kappa^{-1/2}(K_{(1)})\tilde{d}_{2}/(T_{2}(K_{(1)})\geq 2\eta^{\star}(K_{(1)})$ and $\lceil\lceil\log_{c}(\kappa^{1/2}(K_{(1)})\cdot T_{2}(K_{(1)}))\rceil\cdot 4\eta^{\star 2}(K_{(1)})/\sigma^{2}\rceil\leq\tilde{d}_{1}^{2}/(C\sigma)^{2}$ is violated or
$d_{m}(K_{(1)})\lesssim\tilde{d}_{1}/c.$
We note here that $\eta^{\star}(K)\geq\eta^{\star}(K_{(1)})$ since we have the inclusion $K_{(1)}\subseteq K$. Hence, by Lemma 2.9, we can find an estimator $\tilde{\mu}_{2}$, such that
$\|\tilde{\mu}_{2}-(\mu-0)/2\|_{2}\leq\tilde{d}_{1}/L,$
with probability at least $1-\exp(-\tilde{d}_{1}/(\sigma^{2}C^{{}^{\prime}2}))-1/(1+\tilde{d}_{1}^{3}/\sigma^{3})$. Next consider $\bar{\mu}_{2}=\Pi_{K_{(1)}}\tilde{\mu}_{2}$ where we use $\epsilon=\tilde{d}_{1}/L$ in Lemma A.1. We conclude that
$\|\bar{\mu}_{2}-(\mu-0)/2\|_{2}$ $\leq\|\tilde{\mu}_{2}-(\mu-0)/2\|_{2}+\sqrt{\epsilon^{2}+2\epsilon\|\tilde{\mu}_{2}-(\mu-0)/2\|_{2}}$
$\leq(\sqrt{3}+1)\tilde{d}_{1}/L$
$=\tilde{d}_{1}/\tilde{L}$.
Next consider the estimator $\hat{\mu}_{2}:=\Pi_{K}(2\bar{\mu}_{2}+0)$ with $\epsilon=2\tilde{d}_{1}/\tilde{L}$ in Lemma A.1. We have
$\|\Pi_{K}(2\bar{\mu}_{2}+0)-\mu\|_{2}$ $\leq\|2\bar{\mu}_{2}+0-\mu\|_{2}+\sqrt{\epsilon^{2}+2\epsilon\|2\bar{\mu}_{2}+0-\mu\|_{2}}$
$\leq 2(\sqrt{3}+1)\tilde{d}_{1}/\tilde{L}=\tilde{d}_{2}.$
Assume now that $\tilde{d}_{2}\geq 2r\vee C\sigma$ so that we do not output $\hat{\mu}_{2}$. Apply now Lemma 2.1 to the set $K_{(2)}$. Set $\tilde{d}_{2}=2(\sqrt{3}+1)\tilde{d}_{1}/\tilde{L}$. We know that either one of $\kappa^{-1/2}(K_{(2)})\tilde{d}_{2}/(T_{2}(K_{(2)})\geq 2\eta^{\star}(K_{(2)})$ and $\lceil\lceil\log_{c}(\kappa^{1/2}(K_{(2)})\cdot T_{2}(K_{(2)}))\rceil\cdot 4\eta^{\star 2}(K_{(2)})/\sigma^{2}\rceil\leq\tilde{d}_{2}^{2}/(C\sigma)^{2}$ is violated or
$d_{m}(K_{(2)})\lesssim\tilde{d}_{2}/c.$
We note here that $\eta^{\star}(K)\geq\eta^{\star}(K_{(2)})$ since we have the inclusion $K_{(2)}\subseteq K$. Hence by Lemma 2.9, we can find an estimator $\tilde{\mu}_{3}$, such that
$\|\tilde{\mu}_{3}-(\mu-\hat{\mu}_{2})/2\|_{2}\leq\tilde{d}_{2}/L,$
with probability at least $1-\exp(-\tilde{d}_{2}/(\sigma^{2}C^{{}^{\prime}2}))-\exp(-\tilde{d}_{1}/(\sigma^{2}C^{{}^{\prime}2}))-1/(1+\tilde{d}_{1}^{3}/\sigma^{3})-1/(1+\tilde{d}_{2}^{3}/\sigma^{3})$. Next consider $\bar{\mu}_{3}=\Pi_{K_{(2)}}\tilde{\mu}_{3}$ where we use $\epsilon=\tilde{d}_{2}/L$ in Lemma A.1. We conclude that
$\|\bar{\mu}_{3}-(\mu-\hat{\mu}_{2})/2\|_{2}$ $\leq\|\tilde{\mu}_{3}-(\mu-\hat{\mu}_{2})/2\|_{2}+\sqrt{\epsilon^{2}+2\epsilon\|\tilde{\mu}_{3}-(\mu-\hat{\mu}_{2})/2\|_{2}}$
$\leq(\sqrt{3}+1)\tilde{d}_{2}/L$
$=\tilde{d}_{2}/\tilde{L}$.
[…]
$\|\Pi_{K}(2\bar{\mu}_{3}+\hat{\mu}_{2})-\mu\|_{2}$ $\leq\|2\bar{\mu}_{3}+\hat{\mu}_{2}-\mu\|_{2}+\sqrt{\epsilon^{2}+2\epsilon\|2\bar{\mu}_{3}+\hat{\mu}_{2}-\mu\|_{2}}$
$\leq 2(\sqrt{3}+1)\tilde{d}_{2}/\tilde{L}=\tilde{d}_{3}.$
We can continue this logic until one of our three running assumptions is violated, or our estimator goes on a low probability path of Lemma 2.9. Let $k\in[M]$ be minimal so that,
$\|\hat{\mu}_{k}-\mu\|_{2}\leq\tilde{d}_{k}=2R\frac{(2(\sqrt{3}+1))^{k-1}}{\tilde{L}^{k-1}},$
but $\|\hat{\mu}_{k+1}-\mu\|_{2}\geq\tilde{d}_{k+1}$, or $\|\hat{\mu}_{k}-\mu\|_{2}\leq\tilde{d}_{k}$ and we have returned $\hat{\mu}_{k}$.
Recall that the estimator $\hat{\mu}_{k+1}=\Pi_{K}(2\bar{\mu}_{k+1}+\hat{\mu}_{k})$. We thus have (for $\epsilon_{k}=2\tilde{d}_{k}/\tilde{L}$)
$\|\Pi_{K}(2\bar{\mu}_{k+1}+\hat{\mu}_{k})-\mu\|_{2}$ $\leq\|2\bar{\mu}_{k+1}+\hat{\mu}_{k}-\mu\|_{2}+\sqrt{\epsilon^{2}+2\epsilon\|2\bar{\mu}_{k+1}+\hat{\mu}_{k}-\mu\|_{2}}$
$\leq 2\|\bar{\mu}_{k+1}\|_{2}+\|\hat{\mu}_{k}-\mu\|_{2}+\sqrt{\epsilon^{2}+2\epsilon(2\|\bar{\mu}_{k+1}\|_{2}+\|\hat{\mu}_{k}-\mu\|_{2})}$
Since $\bar{\mu}_{k+1}\in B(0,\tilde{d}_{k}/2)$ we have that the above is smaller than
$\|\Pi_{K}(2\bar{\mu}_{k+1}+\hat{\mu}_{k})-\mu\|_{2}\leq\tilde{d}_{k}+\|\hat{\mu}_{k}-\mu\|_{2}+\sqrt{\epsilon_{k}^{2}+2\epsilon_{k}(\tilde{d}_{k}+\|\hat{\mu}_{k}-\mu\|_{2})}$
Let $\|\hat{\mu}_{l}-\mu\|_{2}=a_{l}$. This same logic can be generalized for any $l\geq k+1$ to claim that:
$a_{l}$ $=\|\Pi_{K}(2\bar{\mu}_{l}+\hat{\mu}_{l-1})-\mu\|_{2}\leq\tilde{d}_{l-1}+\|\hat{\mu}_{l-1}-\mu\|_{2}+\sqrt{\epsilon_{l-1}^{2}+2\epsilon_{l-1}(\tilde{d}_{l-1}+a_{l-1})}$
$\leq(1+\sqrt{4/\tilde{L}})\tilde{d}_{l-1}+a_{l-1}+\epsilon_{l-1}+\sqrt{2\epsilon_{l-1}a_{l-1}}$
$=a_{l-1}+(1+\sqrt{4/\tilde{L}}+2/\tilde{L})\tilde{d}_{l-1}+\sqrt{2\epsilon_{l-1}a_{l-1}}$
Next, let $l^{\star}=\mathop{\mathrm{argmax}}_{l\in[k+1,M]}a_{l}$. We note that for $l\in[k+1,M]$ we have $\sqrt{a_{l-1}}\leq\sqrt{a_{l^{\star}}\vee a_{k}}\leq\sqrt{a_{l^{\star}}+a_{k}}\leq\sqrt{a_{l^{\star}}}+\sqrt{\tilde{d}_{k}}$. Then by telescoping the above inequality, we can write
$a_{l^{*}}$ $\leq a_{k}+(1+\sqrt{4/\tilde{L}}+2/\tilde{L})\sum_{i=k+1}^{l^{*}}\tilde{d}_{i-1}+\sum_{i=k+1}^{l^{*}}\sqrt{2\epsilon_{i-1}}(\sqrt{a_{l^{*}}}+\sqrt{\tilde{d}_{k}})$
$\leq a_{k}+(1+\sqrt{4/\tilde{L}}+2/\tilde{L})\sum_{i=k+1}^{M}\tilde{d}_{i-1}+\sum_{i=k+1}^{M}\sqrt{2\epsilon_{i-1}}(\sqrt{a_{l^{*}}}+\sqrt{\tilde{d}_{k}})$
$=\tilde{C}\tilde{d}_{k}+C^{\prime\prime}\sqrt{\tilde{d}_{k}}\sqrt{a_{l^{*}}},$
[…]
$a_{l^{*}}-C^{\prime\prime}\sqrt{\tilde{d}_{k}}\sqrt{a_{l^{*}}}\leq\tilde{C}\tilde{d}_{k}.$
This means that (upon solving the quadratic equation)
$\sqrt{a_{l^{*}}}\leq\frac{C^{\prime\prime}\sqrt{\tilde{d}_{k}}+\sqrt{C^{{}^{\prime}2}\tilde{d}_{k}+4\tilde{C}\tilde{d}_{k}}}{2}\lesssim\sqrt{\tilde{d}_{k}}.$
Thus $\max_{k+1\leq l\leq M}a_{l}\lesssim\tilde{d}_{k}$ which is proportional to the last “correct” value $\|\hat{\mu}_{k}-\mu\|\leq\tilde{d}_{k}$.
There are four reasons why we can “stop” at the $k$-th value. Here by stop we refer to having $\|\hat{\mu}-\mu\|_{2}\leq\tilde{d}_{k}$, but that inequality fails to hold on the next iteration. The first reason for stopping is that our estimator could have gone on a low probability event (we will treat that case shortly).
Second, it is possible that, $\tilde{d}_{k}\leq(2r)\vee(C\sigma)$ and hence we have returned $\hat{\mu}_{k}$. First suppose $2r>C\sigma$. In that case, recall that we are assuming $\sigma\geq r/\sqrt{n}$ (otherwise, we simply output $Y$), implying by Lemma 2.10 that the minimax rate $\eta^{\star}(K)\gtrsim r$. On the other hand, since we are assuming $\|\hat{\mu}_{k}-\mu\|_{2}\lesssim\tilde{d}_{k}$, $\hat{\mu}_{k}$ is rate optimal. Next, suppose $C\sigma\geq 2r$. In that case since $\eta^{\star}(K)\gtrsim(C\sigma)$ (see Prasadan and Neykov (2025b) e.g.) it follows that $\hat{\mu}_{k}$ is rate optimal.
Reasons three and four why we might “halt” a the $k$-th value, are that one of the two running assumptions might be violated, i.e., at least one of $\kappa^{-1/2}(K_{(k)})\tilde{d}_{k}/(T_{2}(K_{(k)}))<2\eta^{\star}(K_{(k)})$ or $\lceil\lceil\log_{c}(\kappa^{1/2}(K_{(k)})\cdot T_{2}(K_{(k)}))\rceil\cdot 4\eta^{\star 2}(K_{(k)})/\sigma^{2}\rceil>\tilde{d}_{k}^{2}/(C\sigma)^{2}$ holds which ensures near optimality upon taking into account the fact that $\eta^{\star}(K_{(k)})\leq\eta^{\star}(K)$.
Using steps similar to the proofs of Lemmas B.5, B.6 and B.7 of Prasadan and Neykov (2024) along with Lemma 2.9, we can show that the accumulated probability bounds for $t\gtrsim\bigg[2\sqrt{2}T_{2}(K)\cdot\max_{k\in[M]}\kappa^{1/2}(K_{(k)})\vee\sqrt{4C^{2}\log_{c}(\max_{k\in[M]}\kappa^{1/2}(K_{(k)})\cdot\sqrt{2}T_{2}(K))}\bigg]\eta^{\star}(K)$, are:
$\mathbb{P}(\|\hat{\mu}-\mu\|_{2}\geq t)\leq\sum_{i:\tilde{d}_{i}\gtrsim t}\exp(-\tilde{d}_{i}^{2}/(\sigma^{2}C^{2}))+\frac{1}{1+\tilde{d}_{i}^{3}/(\sigma^{3})}\leq\exp(-t^{2}/\sigma^{2}\tilde{C}^{2})+\tilde{C}\sigma^{3}/t^{3}.$ (2.11)
We remind the reader that in the inequality above, the bounds $\frac{1}{1+\tilde{d}_{i}^{3}/(\sigma^{3})}$ come from the potential failure of our oracle $\mathcal{O}_{K_{(j)}}(\cdot)$. We remark here that $r\leq\sqrt{n}\sigma$ (otherwise we simply output $Y$ which is optimal), implies that $\log R/\sigma$ is polynomial in $n$, ensuring that all iterations that we make in Section 3 to calculate the oracle $\mathcal{O}_{K_{(j)}}$ are polynomial in $n$ (recall Remark 2.6). Observe that the above bound continues to hold for values of $t\geq d$ as well (the RHS is actually simply 0 then, which is trivially upper bounded by the exponential). Note that for smaller values of $t$ we do not need any probability bounds since we have the trivial:
$\mathbb{P}(\|\hat{\mu}-\mu\|_{2}\geq t)\leq 1.$
Hence after integration we have that
$\mathbb{E}\|\hat{\mu}-\mu\|_{2}^{2}\lesssim\Upsilon(K)\cdot\eta^{\star 2}+\sigma^{2}\exp(-\Upsilon(K)\cdot\eta^{\star 2}/(\tilde{C}^{\prime}\sigma)^{2})+\sigma^{3}/(\sqrt{\Upsilon(K)}\eta^{\star}(K))\lesssim\Upsilon(K)\cdot\eta^{\star 2},$
where $\lesssim$ hides absolute constant factors and $\Upsilon(K)\geq 1$ depends on $T_{2}(K)$ and $\max_{k\in[M]}\kappa(K_{(k)})$ as described above. The final inequality holds true since, $\sigma^{2}\lesssim\eta^{\star 2}(K)$ (as shown in Prasadan and Neykov (2025b) e.g.).
Also if we have outputted $Y$ in the case when $r\geq\sqrt{n}\sigma$ then it is clear that $\mathbb{E}\|Y-\mu\|_{2}^{2}=n\underline{\sigma}^{2}\leq n\sigma^{2}$ which is also optimal by Lemma 2.10.
3 Quadratic Form Maximization (QFM)
In this section, we review some of the results of Bhattiprolu et al. (2021), and in addition, we establish some new results regarding QFM over convex bodies. Coupled with the results from Section 2 this demonstrates that our algorithm works unconditionally for plenty of type-2, well-balanced convex bodies.
The main problem of interest in this seciton is QFM over an origin-symmetric type-2 body $K$. Specifically given a matrix $X$ we want to appxoximately maximize
$\max_{p\in K}p^{\top}Xp.$
Here by approximate maximization, we mean finding a point $\mathcal{O}(X)\in K$ such that $\mathcal{O}(X)^{\top}X\mathcal{O}(X)\geq\kappa^{-1}\max_{p\in K}p^{\top}Xp$, for some $\kappa\geq 1$ which may vary with the set $K$ but is sufficently small.
This is a diffuclt problem, but fortunately the paper Bhattiprolu et al. (2021) studied it in great detail and provided polynomial-time algorithms based on the ellipsoid algorithm for a plethora of origin symmetric type-2 convex bodies. In our case the matrix $X$ is symmetric psd. In Bhattiprolu et al. (2021), the authors considered the same setting, and even the broader setting when $X$ is simply symmetric.
We will now review some of their results. Bhattiprolu et al. propose two relaxations of the problem. Let $g\sim N(0,I)$ be a standard Gaussian vector. They look at the following non-convex but approximately convex problem:
$\mathop{\mathrm{argmax}}_{W}\langle X,W\rangle$
$\mbox{subject to }\mathbb{E}\rho^{2}_{K}(W^{1/2}g)\leq 1$
$~~~~~~~~~~~~~~W\geq 0,W^{\top}=W$
The second problem they formulate is:
$\mathop{\mathrm{argmax}}_{W}\rho_{K^{\circ}}(X^{1/2}W^{1/2}g)$
$\mbox{subject to }\operatorname{tr}(W)\leq 1$
$~~~~~~~~~~~~~~W\geq 0,W^{\top}=W,$
where $K^{\circ}=\{v\in\mathbb{R}^{n}:\sup_{x\in K}\langle v,x\rangle\leq 1\}$ is the polar body of $K$ so that $\rho_{K^{\circ}}$ is the dual norm of $\rho_{K}$, and $X^{1/2}$ is a symmetric square root of $X$ which is symmetric and psd. It is proved that these two problems, have optimal values equal to that of the correpsonding QFM problem. After approximately solving one of the the two problems, rounding algorithms are proposed to obtain the vectors $\mathcal{O}(X)$ from the optimal matrix $W^{\star}(X)$.
In order to solve the two problems, Bhattiprolu et al. (2021) proceed to define the lower and upper covariance regions. The upper covariance region is defined as:
$\mathcal{U}(K):=\{W\text{ symmetric psd}:\mathbb{E}\rho^{2}_{K}(W^{1/2}g)\leq 1\}$
while the lower covariance region is
$\mathcal{L}(K):=\{W\text{ symmetric psd}:\mathbb{E}\rho^{2}_{K^{\circ}}(W^{1/2}g)\geq 1\}.$
Bhattiprolu et al. (2021) observe that
$\mathcal{U}(K)\subseteq\operatorname{conv}(\mathcal{U}(K))\subseteq T_{2}(K)^{2}\cdot\mathcal{U}(K)$
$\mathcal{L}(K)\subseteq\operatorname{conv}(\mathcal{L}(K))\subseteq\frac{1}{C_{2}(K^{\circ})}\cdot\mathcal{L}(K),$
where $C_{2}(K^{\circ})$ is the cotype constant of $K^{\circ}$ which is known to satisfy
$C_{2}(K^{\circ})\leq T_{2}(K)\leq C_{2}(K^{\circ})\log(n+1),$
where, as above, $K^{\circ}$ is the polar body of $K$. In that sense the upper and lower covariance sets are approximately convex. In their sections 3 and 4 the authors of Bhattiprolu et al. (2021) establish conditional polynomial-time algorithms. These algorithms, which are based on an (approximate) ellipsoid algorithms, require the existence of an approximate separation oracle for (at least) one of the two sets $\mathcal{U}(K),\mathcal{L}(K)$. The reader is kindly refered to Definitions 1.6 and 1.11 in Bhattiprolu et al. (2021) for what an approximate separation oracle means. The conditional [^10] Here the word conditional means that these algorithms work assuming separation oracles for one of the upper or lower covariance bodies. maximization polynomial-time algorithms guarantees are given in Propositions 4.2 and 4.4 of Bhattiprolu et al. (2021). Importantly, we note that the algorithms in Propositions 4.2 and 4.4 not only can provide a near optimal value but also have rounding algorithms which produce near optimal vectors with high probability after polynomially many rounds. In addition to the approximate separation oracles for $\mathcal{U}(K)$ or $\mathcal{L}(K)$, these results require the set $K$ to be balanced in the sense that $rB_{2}\subseteq K\subseteq RB_{2}$ for some known constants $R,r\in\mathbb{R}_{+}$. The complexity of the algorithms is polynomial in $n,\log R/r$ and $\operatorname{bit}(X)$. Section 6 of Bhattiprolu et al. (2021) shows that the class of convex bodies for which we can evaluate an approximate separation oracle of the upper covariance region $\mathcal{U}(K)$ is closed under variety of operations including intersection, Minkowski addition, subspaces and quotients. This implies in particular that if one has a separation oracle for $\mathcal{U}(K)$ one is able to maximize over the set $K\cap cB_{2}$ for any $c\geq r$, since it is easy to see [^11] Indeed results in Section 7 of Bhattiprolu et al. (2021) show that any type-2 exactly 2-convex and sign-invariant norm like $\|x\|_{2}/c$ (the Minkowski gauge of $cB_{2}$), admits an approximate separation oracle of $\mathcal{U}(cB_{2})$, and hence by results of Section 6 the intersection set $K\cap cB_{2}$ also admits such an oracle). See also the proof of Theorem 3.3 where we explain why explicit separation oracle of the convex set $\mathcal{U}(cB_{2})$ exists. that the set $\mathcal{U}(cB_{2})$ admits a separation oracle. Next, in Section 7 Bhattiprolu et al. (2021) exhibit unconditional algorithms for a plethora of convex bodies $K$, where one has to have an oracle computing $\rho_{K}(x)$ for any $x\in\mathbb{R}^{n}$ (see Theorem 7.6 and Theorem Theorem 7.12). In other words, the authors are able to construct the approximate separation oracles for the upper and lower covariance regions of these sets. Specifically, the assumptions on the Minkowski gauges of the convex bodies in the latter theorems are as follows:
  • $\rho_{K}$ is a sign-invariant (i.e. $\rho_{K}((x_{1},\ldots,x_{n}))=\rho_{K}((\eta_{j}x_{j})_{j\in[n]})$ for any $x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$ and any $\eta_{j}\in\{\pm 1\}$), and exactly 2-convex norm:
    $\rho_{K}((\sum_{i\in[m]}|x_{i}|^{2})^{1/2})\leq(\sum_{i\in[m]}\rho^{2}_{K}(x_{i}))^{1/2},$
    for any $m\in\mathbb{N}$ and vectors $x_{i}\in\mathbb{R}^{n}$, where $|\cdot|^{2}$ is applied entry-wise.
  • $\rho_{K}$ is a symmetric norm, i.e.,
    $\rho_{K}((x_{1},\ldots,x_{n}))=\rho_{K}((\eta_{j}x_{\pi_{j}})_{j\in[n]}),$
    for any $x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$, any $\eta_{j}\in\{\pm 1\}$ and an arbitrary permutation $\pi$.
If $\rho_{K}$ is symmetric, then one can construct a separation oracle for $\mathcal{L}(K)$, and hence can perform QFM over $K$. Observe also that when $\rho_{K}(\cdot)$ is symmetric, the Minkowski gauge [^12] We have $\rho_{K\cap B(0,c)}(x)=\rho_{K}(x)\vee\|x\|_{2}/c$. of the intersection of $K\cap B(0,c)$ for some $c\in\mathbb{R}^{+}$, is also symmetric. Thus if the original set $K$ is given by a symmetric norm $\rho_{K}$ one can implement the algorithm from Section 2 unconditionally in polynomial-time.
If $\rho_{K}$ is sign-invariant and exaclty 2-convex (such as any $\ell_{p}$ norm with $p\geq 2$), one can find a separation oracle to $\mathcal{U}(K)$. Hence by our comment above, it follows that one can perform QFM over $K\cap cB_{2}$ for any $r\leq c\leq R$ and thus one can implement the algorithm from Section 2 unconditionally in polynomial-time.
In conjunction with Section 2, these results imply that for any such convex bodies we can find the mean vector at a nearly minimax optimal rate. We now expand the class of bodies for which unconditional algorithms exist.
Theorem 3.1.
Let $\|\cdot\|$ be an exactly $2$-convex sign-symmetric norm on $\mathbb{R}^{m}$ [^13] Here $m\in\mathbb{N}$ denotes an arbitrary integer and is unrelated to the parameter $m$ defined in Section 2. for $m\geq n$ which is also a type-2 norm with small constant which we denote by $T_{2}(m)$. Assume that we have an oracle for evaluating $\|\cdot\|$. Suppose our convex body $K=\{x\in\mathbb{R}^{n}:\|Ax\|\leq 1\}$, where we additionally assume that there are known $r,R\in\mathbb{R}_{+}$ such that $rB_{2}\subseteq K\subseteq RB_{2}$. Then for any $r\leq c\leq R$, we can find an approximate maximizer of $\mathop{\mathrm{argmax}}_{p\in K\cap cB_{2}}p^{\top}Xp$: $\mathcal{O}(X)$ for any psd matrix $X$ in polynomial-time operations: $\operatorname{poly}(m,\log 1/q,\log R/r,\operatorname{bit}(X),\operatorname{bit}(A))$ with probability at least 1q1-q. Here $\kappa(K\cap cB_{2})\leq 8eT_{2}^{2}(K)\log m\leq 8eT_{2}^{2}(m)\log m$.
Proof.
We first establish that we can approximately optimize $\mathop{\mathrm{argmax}}_{p\in K}p^{\top}Xp$.
Using the first relaxation given by Bhattiprolu et al. (2021) we can rewrite the problem as
$\mathop{\mathrm{argmax}}\langle X,W\rangle$ (3.1)
$\mbox{subject to }\mathbb{E}\rho_{K}^{2}(W^{1/2}g)=\mathbb{E}\|(AWA^{\top})^{1/2}g^{\prime}\|^{2}\leq 1$
$~~~~~~~~~~~~~~W\geq 0,W=W^{\top},W\in\mathbb{R}^{m\times m}$
Above $g\in\mathbb{R}^{n},g^{\prime}\in\mathbb{R}^{m}$ are stanfard Gaussian vectors and $(AXA^{\top})$ is a $n\times n$ matrix.
Lemma 3.2.
Program (3.1) has a value which coincides with the QFM over the set $\max_{p\in K}p^{\top}Xp$. Let $Q_{\max}(X)$ be this value.
Proof.
The proof is identical to Observation 4.1 of Bhattiprolu et al. (2021). We have
$\langle X,W\rangle=\mathbb{E}\langle W^{1/2}g,XW^{1/2}g\rangle\leq Q_{\max}(X)\cdot\mathbb{E}(\rho_{K}^{2}(W^{1/2}g)),$
where $g\sim N(0,I_{n})$. Now by assumption $\rho_{K}(p)=\|Ap\|$ for some type-2, exactly 2-convex norm $\|\cdot\|$. Thus
$\rho_{K}^{2}(W^{1/2}g)=\|AW^{1/2}g\|^{2}\stackrel{{\scriptstyle d}}{{=}}\|(AWA^{\top})^{1/2}g^{\prime}\|^{2},$
for some gaussian vector $g^{\prime}\sim N(0,I_{m})$. The latter identity comes from the fact that $AW^{1/2}g$ is a Gaussian vector with covariance $AXA^{\top}$. This implies $\langle X,W\rangle\leq Q_{\max}(X)$.
Finally note that we can take $W=ww^{\top}$ where $w\in K$ achieves the $Q_{\max}(X)$. We have that $W^{1/2}=ww^{\top}/\|w\|_{2}$ and thus, $\mathbb{E}\rho_{K}^{2}(W^{1/2}g)=\rho_{K}^{2}(w)\mathbb{E}(w^{\top}g)^{2}/\|w\|_{2}=1$, so that $W$ is a feasible matrix which attains the value $Q_{\max}(X)$. ∎
Define the function
$f(W)=\|\operatorname{diag}(AWA^{\top})^{1/2}\|^{2},$
where the 1/2 is taken entry-wise on the non-negative vector $\operatorname{diag}(AWA^{\top})$. This is a convex function according to the proof of Theorem 7.6 of Bhattiprolu et al. (2021). Specifically, the proof of Theorem 7.6 argues that the map $Y\mapsto\|\operatorname{diag}(Y)^{1/2}\|^{2}$ is convex for any psd matrix $Y\in\mathbb{R}^{m\times m}$. Since the map $Y\mapsto AYA^{\top}$ is affine: $A(\theta Y_{1}+(1-\theta)Y_{2})A^{\top}=\theta AY_{1}A^{\top}+(1-\theta)AY_{2}A^{\top}$, then it follows that the map $W\mapsto\|\operatorname{diag}(AWA^{\top})^{1/2}\|^{2}$ is also convex.
Using the proof of Theorem 7.6 (and Theorem 7.4) of Bhattiprolu et al. (2021) and the properties of $\|\cdot\|$, we can claim that
$f(W)\leq\mathbb{E}\rho^{2}_{K}(W^{1/2}g)\leq\beta f(W),$
where $\beta\leq(1+o(1))(e\log m)$. Next, we can apply Corollary 4.12 part 1. to claim that we can find an $\beta\cdot T^{2}_{2}(K)$ approximate maximizer to program (3.1) in polynomial-time. Finally, once we find a matrix $W$ for this problem, we need to output a vector $\mathcal{O}(X)$.
We know by the properties of $W$ that $\langle X,W\rangle=\mathbb{E}\langle W^{1/2}g,XW^{1/2}g\rangle\geq(1-o(1))\operatorname{OPT}/(\beta\cdot T^{2}_{2}(K))$, where $\operatorname{OPT}$ is the optimal value of program (3.1). Next, following the idea of the rounding algorithm from Proposition 4.2 of Bhattiprolu et al. (2021) we generate Gaussian variables $g\in\mathbb{R}^{n}$ and compute, $G:=\langle W^{1/2}g,XW^{1/2}g\rangle-(1-o(1))\|AW^{1/2}g\|^{2}\operatorname{OPT}/((1+\varepsilon)\beta\cdot T^{2}_{2}(K))$, for some small $\epsilon>0$, until the number is positive for one of the vectors $g$. By Chebyshev’s inequality we can show that this will happen in polynomially many steps. Next, the approximately optimal solution is $W^{1/2}g/\|AW^{1/2}g\|$. We will now explicitly show the argument with Chebyshev’s inequality. Taking expectation we have
$\mathbb{E}G$ $\geq(1-o(1))\operatorname{OPT}/(\beta\cdot T_{2}^{2}(K))-(1-o(1))\operatorname{OPT}/((1+\varepsilon)\beta\cdot T^{2}_{2}(K))$
$\geq(1-o(1))\operatorname{OPT}\epsilon/((1+\epsilon)\beta\cdot T_{2}^{2}(K))\geq 0,$
where we used the fact that $\mathbb{E}\|AW^{1/2}g\|^{2}\leq 1$ by assumption. Now,
$\operatorname{Var}(\sum_{i\in[t]}G_{i})$ $\leq t\mathbb{E}[G^{2}]\leq 2t[\mathbb{E}\langle W^{1/2}g,XW^{1/2}g\rangle^{2}+\mathbb{E}\|AW^{1/2}g\|^{4}\operatorname{OPT}^{2}/((1+\epsilon)\beta\cdot T_{2}^{2}(K))^{2}]$
$=2t\mathbb{E}\|X^{1/2}W^{1/2}g\|_{2}^{4}+2t\mathbb{E}\|AW^{1/2}g\|^{4}\operatorname{OPT}^{2}/((1+\epsilon)\beta\cdot T_{2}^{2}(K))^{2}]$
$\lesssim t(\mathbb{E}\|X^{1/2}W^{1/2}g\|_{2}^{2})^{2}+t(\mathbb{E}\|AW^{1/2}g\|^{2})^{2}\operatorname{OPT}^{2}/((1+\epsilon)\beta\cdot T_{2}^{2}(K))^{2}]$
$\leq t\operatorname{OPT}^{2}+t\operatorname{OPT}^{2}/((1+\epsilon)\beta\cdot T_{2}^{2}(K))^{2}]$
$=t\operatorname{OPT}^{2}\delta,$
where the penultimate inequality follows from Khintchine-Kahane’s theorem (see Theorem 2.1 in Bhattiprolu et al. (2021) e.g.), and $\delta:=1+\frac{1}{(1+\epsilon)\beta\cdot T_{2}^{2}(K))^{2}}$. Thus by Chebyshev’s inequality with probability at least $1-1/k^{2}$:
$\max_{i\in[t]}G_{i}\geq\sum_{i\in[t]}G_{i}/t\geq\mathbb{E}G-kt^{-1/2}\sqrt{\delta}\operatorname{OPT}.$
Thus if $k$ is of constant order, and $t\gtrsim\epsilon^{-2}$ we will have the above will be positive. The probability can be boosted to 1q1-q for some $q\in[0,1]$, at the expense of generating $tk^{2}\log q^{-1}$ Gaussians and taking the maximum. This is because if we run the above procedure $i$ times, we will obtain a bound $(1/k^{2})^{i}\leq\exp(-i/k^{2})$ to see at least one positive variable.
Now we turn to the problem where we need to find the optimum for the intersection set $K\cap cB_{2}$. This set is clearly balanced since $r\leq c\leq R$. Furthermore, we have
$\mathbb{E}\rho^{2}_{K}(W^{1/2}g)\vee\mathbb{E}\|W^{1/2}g\|^{2}_{2}\cdot c^{-2}$
$\leq\mathbb{E}[\rho^{2}_{K}(W^{1/2}g)\vee\|W^{1/2}g\|^{2}_{2}\cdot c^{-2}]$
$\leq 2(\mathbb{E}\rho^{2}_{K}(W^{1/2}g)\vee\mathbb{E}\|W^{1/2}g\|^{2}_{2}\cdot c^{-2})$
Since $\mathbb{E}\|W^{1/2}g\|_{2}^{2}=\operatorname{tr}(W)$ and we know that $f(W)\leq\mathbb{E}\rho^{2}_{K}(W^{1/2}g)\leq\beta f(W)$, the convex function $f(W)\vee[\operatorname{tr}(W)c^{-2}]$ satisfies
$f(W)\vee[\operatorname{tr}(W)c^{-2}]\leq\mathbb{E}[\rho^{2}_{K}(W^{1/2}g)\vee\|W^{1/2}g\|^{2}_{2}\cdot c^{-2}]\leq 2\beta(f(W)\vee[\operatorname{tr}(W)c^{-2}])$
Thus again by Corollary 4.12 part 1. we can claim, that we can find a matrix $W$ which is at most $2\beta\cdot T^{2}_{2}(K\cap cB_{2})\leq 4\beta\cdot\max(T^{2}_{2}(K),1)=4\beta\cdot T^{2}_{2}(K)$ approximate maximizer. Now to the rounding algorithm. The logic is the same as before: we know that $W$ that $\mathbb{E}\langle W^{1/2}g,XW^{1/2}g\rangle\geq(1-o(1))\operatorname{OPT}/(\beta\cdot T^{2}_{2}(K))$. We then generate polynomially many Gaussian variables $g\in\mathbb{R}^{n}$ and compute, $\langle W^{1/2}g,CW^{1/2}g\rangle-(1-o(1))(\rho_{K}(W^{1/2}g)\vee[\|W^{1/2}g\|_{2}c^{-1}])^{2}\operatorname{OPT}/((1+\varepsilon)\beta\cdot T^{2}_{2}(K))$. And the maximum $g$ gives a $W^{1/2}g/(\rho_{K}(W^{1/2}g)\vee[\|W^{1/2}g\|_{2}c^{-1}])$, which is our approximate solution. As before the probability can be boosted to 1q1-q with $c\log q^{-1}$ generations of Gaussians. ∎
Theorem 3.3.
Assume we have a symmetric type-2 norm $\|\cdot\|$ and an oracle for evaluation of $\|\cdot\|$. Denote the type-2 constant of $\|\cdot\|$ with $T_{2}(n)$. Let the Minkowski gauge of the body $K$ be given by $\rho_{K}(x)=\|Ax\|$ where $A_{n\times n}$ is a given full rank matrix. Suppose that the set $K$ is balanced in the sense that there are known $(r,R)$ so that $rB_{2}\subseteq K\subseteq RB_{2}$. Then for any $r\leq c\leq R$, we can find an approximate maximizer of $\mathop{\mathrm{argmax}}_{p\in K\cap cB_{2}}p^{\top}Xp$: $\mathcal{O}(X)$ for any psd matrix $X$ in polynomial-time operations: $\operatorname{poly}(n,\log 1/q,\log R/r,\operatorname{bit}(X),\operatorname{bit}(A))$ with probability at least 1q1-q. Here $\kappa(K\cap cB_{2})\lesssim 2T_{2}^{10}(K)\log^{4}(eT_{2}(K))\leq 2T_{2}^{10}(n)\log^{4}(eT_{2}(n))$.
Proof.
For any input psd symmetric matrix $X$ we have the following identity:
$\max_{p\in K}p^{\top}Xp=\max_{p:\|Ap\|\leq 1}p^{\top}Xp=\max_{x:\|x\|\leq 1}x^{\top}A^{-1\top}XA^{-1}x.$
Thus, we can approximately optimize the above problem by the results of Section 7 of Bhattiprolu et al. (2021), see Theorem 7.12. Here $\kappa\lesssim T_{2}^{6}(K)\log^{4}(eT_{2}(K))$.
Using Proposition 4.13 of Bhattiprolu et al. (2021) we can now claim that we have a $\alpha\lesssim T_{2}^{8}(K)\log^{4}(eT_{2}(K))$-approximate separation oracle for the set $\mathcal{U}(K)$ where $K=\{x\in\mathbb{R}^{n}:\|Ax\|\leq 1\}$. Next, using Proposition 6.5 of Bhattiprolu et al. (2021) we know we can obtain a polynomial-time approximate quadratic maximization algorithm over the set $K\cap cB_{2}$ for any $r\leq c\leq R$. Not only that but we will also have rounding algorithms to obtain $\mathcal{O}(X)$ (we can simply replicate the rounding algorithms logic from the proof of Theorem 3.1 above which describes how to obtain rounding algorithms with the probability guarantees that we require). Here the constant $\kappa\lesssim 2\max(\alpha,1)\cdot\max(T^{2}_{2}(K),1)$.
In the above we used the fact that we can find an exact separation oracle for $\mathcal{U}(cB_{2})$. This is easy to see since $\mathcal{U}(cB_{2})=\{W\geq 0,W=W^{\top}:\operatorname{tr}(W)\leq c\}$ is convex and admits a simple explicit separation oracle.
4 Robust Heavy-Tailed Extension
Consider the data $\tilde{Y}_{i}=\mu+\xi_{i}$, $i\in[N]$ where $\xi_{i}$ are i.i.d. vectors on $\mathbb{R}^{n}$ such that $\mathbb{E}\xi_{i}\xi_{i}^{\top}=\Sigma$ with $\|\Sigma\|_{\operatorname{op}}\leq\sigma^{2}$ where $\sigma^{2}$ (or an upper bound) is known to the statistician. Next the data is manipulated by a malicious and all-powerful adversary so that a (known) fraction $\varepsilon\leq c_{0}<1/2$ of the observations can be arbitrary. This adversary is allowed to inspect all samples first, pick the ones to switch, and is also given oracle knowledge to all parameters of the model including $\mu$, $K$, $\sigma$ and oracle knowledge of our algorithm. We denote the observations after the adversary has tampered with the data as $(Y_{i})_{i\in[N]}$.
We are interested in estimating $\mu$ in a minimax optimal way. In order to derive the rate Peng et al. (2025), suppose that $N\gtrsim\sup_{\delta>0}\log M^{\operatorname{loc}}_{K}(\delta)$. Since $n\gtrsim\log M^{\operatorname{loc}}_{K}(\delta)$ this assumption is implied if $N\gtrsim n$. On the other hand, since we assume that the set $K$ is well-balanced, we have that $rB_{2}\subseteq K$ and therefore, when $\delta\lesssim r$, $\log M^{\operatorname{loc}}_{K}(\delta)\gtrsim n$. Thus, in this section, we will assume the assumption: $N\gtrsim n$. In that case, as shown by Peng et al. (2025) the minimax rate $\inf_{\hat{\mu}}\sup_{\mu\in K}\sup_{\mathcal{C}}\mathbb{E}\|\hat{\mu}-\mu\|_{2}^{2}$ is $\asymp\eta^{\star 2}\wedge d^{2}$ where $\eta^{\star}=(\delta^{\star}\vee\sqrt{\varepsilon}\sigma)$ where $\delta^{\star}:=\sup\{\delta\geq 0:N\delta^{2}/\sigma^{2}\leq\log M^{\operatorname{loc}}_{K}(\delta)\}$, and where the inner most supremum over $\mathcal{C}$ indicates taking supremum over all possible adversarial schemes.
By substituting the $\sigma$ with $\sigma/\sqrt{N}$ in Lemma 2.1 we obtain the following corollary.
Corollary 4.1.
Let $K$ be a type-2 convex body with type-2 constant $T_{2}(K)$, such that $rB_{2}\subseteq K\subseteq RB_{2}$ for some known $r,R$ with $R=O(1)$. Let $\kappa:=\kappa(K)$ be a given scalar which can depend on $K$. Let $\delta^{\star}:=\delta^{\star}(K)$ be defined through the equation $\delta^{\star}=\sup_{\delta}\{\delta\geq 0:N\delta^{2}/\sigma^{2}\leq\log M_{K}^{\operatorname{loc}}(\delta)\}$. Assume that $\kappa^{-1/2}\tilde{d}/T_{2}(K)\geq 2\delta^{\star}$ and $\lceil\lceil\log_{c}(\kappa^{1/2}\cdot T_{2}(K))\rceil\cdot 4\delta^{\star 2}\cdot N/\sigma^{2}\rceil\leq N\cdot\tilde{d}^{2}/(\sigma^{2}C^{2})$ for $\tilde{d}=2R\geq d$ and where $C>0$ is some absolute constant. Then for $m=\lceil N\cdot\tilde{d}^{2}/(\sigma^{2}C^{2})\rceil\wedge n$, we have
$d_{m}(K)\lesssim\kappa^{-1/2}\cdot\tilde{d}/c,$
We will modify several steps in our algorithm. Specifically, after obtaining the matrix $A^{\star\star}$ for the set $K^{\prime}=K\cap B_{2}(0,\tilde{d}/2)$ we transform the data to $(A^{\star\star}Y_{i})_{i\in[N]}$ and we apply the polynomial-time algorithm of Depersin and Lecué (2022) to the data with $k=u=\lceil N\tilde{d}^{2}/(C^{2}\sigma^{2})\rceil\wedge N$ blocks. In addition, when we obtain the matrix $A^{\star\star}$ we use $N^{-1/2}\sigma$ in place of $\sigma$ in Theorem 2.5 and Algorithm 2.
Now we prove the following lemma
Lemma 4.2.
The estimator $f((A^{\star\star}Y_{i})_{i\in[N]})$, where $f$ denotes the procedure of Depersin and Lecué (2022) with $k=u=\lceil N\tilde{d}^{2}/(C^{2}\sigma^{2})\rceil\wedge N$, of $A^{\star\star}\mu$ satisfies
$\|f((A^{\star\star}Y_{i})_{i\in[N]})-A^{\star\star}\mu\|_{2}^{2}\leq\tilde{d}^{2}/T^{2},$
with probability at least $1-\exp(-k/C^{\prime})-(1+N^{3/2}\tilde{d}^{3}/\sigma^{3})^{-1}$ provided that $k\gtrsim\varepsilon N$, and $\tilde{d}\geq C\sigma/\sqrt{N}$.
Proof.
First, during the estimation of $A^{\star\star}$ our QFM oracle fails with probability at most $1-1/(1+N^{3/2}\tilde{d}^{3}/\sigma^{3})$. Assuming the oracle never fails we now show the result and use the union bound in the end.
From $\tilde{d}\geq C\sigma/\sqrt{N}$ we have that $\lceil N\tilde{d}^{2}/(C^{2}\sigma^{2})\rceil\asymp N\tilde{d}^{2}/(C^{2}\sigma^{2})$. Since $A^{\star\star}\tilde{Y}=A^{\star\star}\mu+A^{\star\star}\xi$, we will first evaluate $\operatorname{tr}(\operatorname{Var}(A^{\star\star}\xi))$ and $\|\operatorname{Var}(A^{\star\star}\xi)\|_{\operatorname{op}}$ as they appear on the right hand side of the main inequality in Theorem 2 of Depersin and Lecué (2022).
For the first term, we see that
$\operatorname{tr}(\operatorname{Var}(A^{\star\star}\xi))$ $=\operatorname{tr}(\mathbb{E}A^{\star\star}\xi\xi^{\top}A^{\star\star})=\operatorname{tr}(\Sigma A^{\star\star 2})$
$\leq\|\Sigma\|_{\operatorname{op}}\|A^{\star\star 2}\|_{*}\leq\|\Sigma\|_{\operatorname{op}}\sum_{i\in[n]}(1-\lambda_{i})=\|\Sigma\|_{\operatorname{op}}m$
$\lesssim N\|\Sigma\|_{\operatorname{op}}\tilde{d}^{2}/(\sigma^{2}C^{2})\leq N\tilde{d}^{2}/C^{2},$
where we remind the reader that $\|\Sigma\|_{\operatorname{op}}\leq\sigma^{2}$. Furthermore, it is clear that
$\|\operatorname{Var}(A^{\star\star}\xi)\|_{\operatorname{op}}=\|\mathbb{E}A^{\star\star}\xi\xi^{\top}A^{\star\star}\|_{\operatorname{op}}\leq\|\Sigma\|_{\operatorname{op}}.$
This, in conjunction with the result of Depersin and Lecué (2022), completes the proof, after the use of the union bound. ∎
Clearly, we can now prove an analogous result to Lemma 2.9 for this setting, with $A_{j}^{\star\star}Y$ substituted with $f((A_{j}^{\star\star}Y_{i})_{i\in[N]})$ with $k=u$ as specified in Lemma 4.2. For the sake of brevity we omit stating this lemma, noting that we denote the analogous constant as $L$. In addition, using Theorem 2 of Depersin and Lecué (2022) with $k=u=N$, on the untransformed data $(Y_{i})_{i\in[N]}$, provided that $\varepsilon$ is smaller than an absolute constant $c_{0}$, with probability at least $1-\exp(-N/C_{0})$ we have
$\|f((Y_{i})_{i\in[N]})-\mu\|_{2}\lesssim\sqrt{n/N}\sigma+\sigma\lesssim\sigma.$ (4.1)
Thus, instead of starting at the 0 vector, we can start at $\hat{\mu}_{0}:=\Pi_{K}f((Y_{i})_{i\in[N]})$, and set $\tilde{d}_{1}=2R\wedge C^{\prime}\sigma$ for some absolute constant $C^{\prime}>0$, and $K_{(1)}=K\cap B_{2}(0,\tilde{d}_{1}/2)$. Here when we write $\Pi_{K}f((Y_{i})_{i\in[N]})$, we mean weak projection as usual. By (4.1) and Lemma A.1 with $\epsilon=\sigma$, we have
$\|\Pi_{K}f((Y_{i})_{i\in[N]})-\mu\|_{2}\lesssim\sigma,$
while $\Pi_{K}f((Y_{i})_{i\in[N]})\in K$. Since we are setting $\epsilon=\sigma$ [^14] Note the distinction between $\varepsilon$ and $\epsilon$: the former is the fraction of outliers, while the latter is simply the parameter of Lemma A.1 in Lemma A.1 this means $\sigma$ cannot be too small. Specifically, we will assume that $\log(1+1/\sigma)$ is polynomial in $n$ and $N$.
The updated algorithm is given in the following:
[…]
4 Near-optimal Robust Mean Estimation Algorithm
0: Convex body $K$, $r,R\in\mathbb{R}_{+}$ such that $rB_{2}\subseteq K\subseteq RB_{2}$, observations $(Y_{i})_{i\in[N]}$, constant $\tilde{L}=L/(\sqrt{3}+1)$ where $L$ is from the robust version of Lemma 2.9; We assume $\tilde{L}>2(\sqrt{3}+1)$
0: Estimator $\hat{\mu}$ of $\mu$ which belongs to the set $K$
1: Overwrite $r:=((\sqrt{n/N}\sigma\vee\sqrt{\varepsilon}\sigma)\wedge r$
2: Set $\tilde{d}_{1}=2R\wedge(C^{\prime}\sigma)$, $K_{(1)}:=K\cap B_{2}(0,\tilde{d}_{1}/2)$, set number of iterations $M=\lceil\log_{\tilde{L}/(2(\sqrt{3}+1))}\tilde{d}_{1}/(2r)\rceil$
3: Set $\hat{\mu}_{1}:=\Pi_{K}f((Y_{i})_{i\in[N]})$ with $\epsilon=\sigma$ in Lemma A.1 and $k=u=N$ (here $f$ denotes the procedure of Depersin and Lecué (2022))
4:for j = 1, …, M do
5:  Find $m=\lceil N\tilde{d}_{j}^{2}/(C^{2}\sigma^{2})\rceil\wedge n$, and calculate the matrix $A_{j}^{\star\star}$ for the set $K_{(j)}$
6:  Find $\tilde{\mu}_{j+1}=(f((A_{j}^{\star\star}Y_{i})_{i\in[N]})-A_{j}^{\star\star}\hat{\mu}_{j})/2$ with $k=u=\lceil N\tilde{d}_{j}^{2}/(C^{2}\sigma^{2})\rceil\wedge N$ for some large constant $C>C^{\prime}$ (here $f$ denotes the procedure of Depersin and Lecué (2022))
7:  Set $\bar{\mu}_{j+1}=\Pi_{K_{(j)}}\tilde{\mu}_{j+1}$ [^15] Here we set $\epsilon=\tilde{d}_{j}/L$ in Lemma A.1 when evaluating the weak projection.
8:  Lift back to $K$: $\hat{\mu}_{j+1}=\Pi_{K}(2\bar{\mu}_{j+1}+\hat{\mu}_{j})$ [^16] Here we set $\epsilon=2\tilde{d}_{j}/\tilde{L}$ in Lemma A.1 when evaluating the weak projection.
9:  Set $\tilde{d}_{j+1}=2(\sqrt{3}+1)\tilde{d}_{j}/\tilde{L}$
10:  Set $K_{(j+1)}:=K\cap B_{2}(0,\tilde{d}_{j+1}/2)$
11:  if $\tilde{d}_{j+1}\leq(2r)\vee(C\sigma/\sqrt{N})$ then
12:   return $\hat{\mu}:=\hat{\mu}_{j+1}$
13:  end if
14:end for
15:return $\hat{\mu}:=\hat{\mu}_{M+1}$
The main theorem of this section resembles the main result of Section 2 with slight modifications.
Theorem 4.3.
Suppose that $T_{2}(K)$ and $\max_{i\in[M]}\kappa(K_{(k)})$ are “small” factors. Assume also that $N\gtrsim n$, and $\log\tilde{d}_{1}/r$ (with the overwritten value of $r$) and $\log(1+1/\sigma)$ are polynomial in $n,N$. Furthermore, let $\frac{d^{2}}{\sigma^{2}}=\frac{\mathrm{diam}(K)^{2}}{\sigma^{2}}\lesssim\exp(N/C_{0})/N$. Then there exists a polynomial up to logarithmic factors function $\Upsilon(K)=g(T_{2}(K),\max_{i\in[M]}\kappa(K_{(k)}))$ such that the output $\hat{\mu}$ of Algorithm 2 satisfies
$\mathbb{E}\|\hat{\mu}-\mu\|_{2}^{2}\lesssim\Upsilon(K)\eta^{\star 2}(K)\wedge d^{2},$
and is thus nearly minimax optimal.
The proof of this result follows the same strategy as the proof of Theorem 2.11, hence we omit it. We now provide a quick sketch of the main changes required. Since here we are assuming $\sigma$ is relatively large, i.e., $\log(1+1/\sigma)$ is polynomial in $n$ and $N$ we can afford to set $r$ to a smaller value in Algorithm 4. The main difference in the two proofs stems from the fact that in the robust case, one needs to assume $k/N=\lceil\tilde{d}_{j}^{2}/(C^{2}\sigma^{2})\rceil\wedge 1\gtrsim\epsilon$, but this condition can be handled as in the proof of Theorem 2.11. This is also precisely the reason why we have a maximum with $\sqrt{\varepsilon}\sigma$ in the definition of $\eta^{\star}$. Also, observe that since for our pilot estimator, $\tilde{d}_{1}\lesssim\sigma$, the number of blocks $k$ will be of the order of $N\tilde{d}_{j}^{2}/(C^{2}\sigma^{2})$ almost always, which implies that the high probability guarantees will be of the form $1-\exp(-N\tilde{d}_{j}^{2}/(C^{\prime}\sigma^{2}))-1/(1+N^{3/2}\tilde{d}_{j}^{3}/\sigma^{3})$ for some $C^{\prime}$. The only exception for this is that the probability bound is capped by $1-\exp(-N/C_{0})$ when $\tilde{d}_{j}\gtrsim\sigma$, by the result of Depersin and Lecué (2022) (observe that here we do not have the term $1/(1+N^{3/2}\tilde{d}_{j}^{3}/\sigma^{3})$ since this term indicates the failure probability for the matrix $A_{j}^{\star\star}$, and if that happens by the logic of the proof of Theorem 2.11, the estimator will have a rate proportional to that of the pilot estimator which is $\lesssim\sigma$ with probability at least $1-\exp(-N/C_{0})$). But since our estimator is proper even if $\|\hat{\mu}-\mu\|_{2}^{2}\gtrsim\sigma^{2}$ we always have the bound $\|\hat{\mu}-\mu\|_{2}^{2}\leq d^{2}$. Hence, when we change the high probability bounds part (2.11), and the integration afterwards, we will have an extra term which is of the form $d^{2}\exp(-N/C_{0})$. Thus if $d^{2}\exp(-N/C_{0})\lesssim\sigma^{2}/N\lesssim\eta^{\star 2}$ we can fold that term in the minimax rate, which is why we assumed it in the theorem. It is also worth noting that the minimax rate on the squared level over the ball $B_{2}(0,r)$ (where here we use the value of $r$ before we overwrite it) is $(n/N\sigma^{2}\vee\epsilon\sigma^{2})\wedge r^{2}$, by a result of Peng et al. (2025). Hence, once we reach $\tilde{d}_{j}\leq 2r$ (with the latter being the overwritten value of $r$) it follows that $\tilde{d}_{j}$ is below the minimax rate of the set $K$.
5 Linear Regression
In this section we consider the problem of linear regression with random design. Suppose we are given $N$ i.i.d. observations from the model:
$Y_{i}=Z_{i}^{\top}\beta+\xi_{i},$
where $\beta\in K\subseteq\mathbb{R}^{n}$, is a type-2 set that satisfies the restrictions of the previous sections. We assume a “moderate dimensional setting”: $N\gtrsim n$ for a some absolute constant. We suppose that the covariates $Z_{i}$ are independent from the noise $\xi_{i}$. We also assume that the diameter of $K$ is bounded above by an absolute constant, i.e., $d:=\mathrm{diam}(K)=O(1)$. Since the diameter is known to be $O(1)$ we will also suppose that the known constant $R=O(1)$, where $rB_{2}\subseteq K\subseteq RB_{2}$. Further, let $\xi_{i}$ be a sub-Gaussian variable with sub-Gaussian parameter $\gamma$ of constant order, i.e., $\mathbb{E}\exp(\lambda\cdot\xi_{i})\leq\exp(\lambda^{2}\gamma^{2}/2)$, for all $\lambda\in\mathbb{R}$ where $\gamma=O(1)$.
Suppose the covariates $Z_{i}$ are centered [^17] If the predictors $Z_{i}$ are not centered one can consider $(Y_{2i}-Y_{2i-1},Z_{2i}-Z_{2i-1})_{i\in[\lfloor N/2\rfloor]}$ to center the predictors while leave the $\beta$ unchanged. We note that this operation prevents the model from having an intercept. sub-Gaussian variables with a well conditioned covariance matrix $\mathbb{E}Z_{i}Z_{i}^{\top}=\Sigma$, i.e., $c^{\prime}\leq\lambda_{\min}(\Sigma)\leq\lambda_{\max}(\Sigma)\leq C^{\prime}$ for some $c^{\prime},C^{\prime}>0$, and sub-Gaussian parameter of constant order, i.e. $\sup_{v\in S^{n-1}}\mathbb{E}\exp(\lambda v^{\top}Z_{i})\leq\exp(\lambda^{2}\zeta^{2}/2)$ for $\zeta=O(1)$. Then the minimax rate $\eta^{\star 2}$, in the case when $\xi_{i}\sim N(0,1)$, is given by the entropy equation $\eta^{\star}=\sup_{\eta}\{\eta\geq 0:N\eta^{2}\leq\log M_{K}^{\operatorname{loc}}(\eta)\}$, as shown in Prasadan and Neykov (2025a).
Returning to Lemma 2.1, and plugging in $N^{-1/2}$ for $\sigma$ we obtain the following corollary.
Corollary 5.1.
Let $K$ be a type-2 convex body with type-2 constant $T_{2}(K)$, such that $rB_{2}\subseteq K\subseteq RB_{2}$ for some known $r,R$ with $R=O(1)$. Let $\kappa:=\kappa(K)$ be a given scalar which can depend on $K$. Let $\eta^{\star}:=\eta^{\star}(K)$ be the minimax rate defined by equation $\eta^{\star}=\sup_{\eta}\{\eta\geq 0:N\eta^{2}\leq\log M_{K}^{\operatorname{loc}}(\eta)\}$. Assume that $\kappa^{-1/2}\tilde{d}/T_{2}(K)\geq 2\eta^{\star}$ and $\lceil\lceil\log_{c}(\kappa^{1/2}\cdot T_{2}(K))\rceil\cdot 4\eta^{\star 2}\cdot N\rceil\leq N\cdot\tilde{d}^{2}/C^{2}$ for $\tilde{d}=2R\geq d$ and where $C>0$ is some absolute constant. Then for $m=\lceil N\cdot\tilde{d}^{2}/C^{2}\rceil\wedge n$, we have
$d_{m}(K)\lesssim\kappa^{-1/2}\cdot\tilde{d}/c,$
We now prove a lemma analogous to Lemma 2.9. However, the proof is more involved and uses an estimator adapted to the regression setting. We also mention that we are using Theorem 2.5 and Algorithm 2 with $N^{-1/2}$ for $\sigma$.
Lemma 5.2.
Suppose that we have an estimator $\hat{\beta}_{j}$ of $\beta$ such that $\|\hat{\beta}_{j}-\beta\|_{2}\leq\tilde{d}_{j}$ happens with high probability, for some $\tilde{d}_{j}\in\mathbb{R}_{+}$ such that $\tilde{d}_{j}\geq(2r)\vee(C/\sqrt{N})$. Assume also $N\gtrsim n$ for some absolute constant. Let $K_{(j)}=K\cap B(0,\tilde{d}_{j}/2)$. Obtain $m$ and the matrix $A_{j}^{\star\star}$ from Corollary 5.1 and Algorithm 2 for the set $K_{(j)}$. Let
$\beta^{\prime}_{j+1}:=\mathop{\mathrm{argmin}}_{\nu\in K_{(j)}}\sum_{i\in[N]}((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})/2-Z_{i}^{\top}A_{j}^{\star\star}\nu)^{2}$ (5.1)
and define the estimator $\tilde{\beta}_{j+1}:=A^{\star\star}_{j}\beta^{\prime}_{j+1}$. Then solving program (5.1) can be approximated in polynomial-time and the estimator $\tilde{\beta}_{j+1}$ satisfies:
$\mathbb{P}(\|\tilde{\beta}_{j+1}-(\beta-\hat{\beta}_{j})/2\|_{2}\geq\tilde{d}_{j}/L,\|\hat{\beta}_{j}-\beta\|_{2}\leq\tilde{d}_{j})\leq n\exp(-Nc^{\prime}\tilde{d}_{j}^{2})+(1+N^{3/2}\tilde{d}_{j}^{3})^{-1}.$
for some absolute constant $L$ which can be made large provided that the constant $C$ is selected to be sufficiently large.
Proof.
The QFM oracle will make a mistake in Algorithm 2 with probability at most $1/(1+N^{3/2}\tilde{d}_{j}^{3})$. We now place ourselves on the event that the matrix $A_{j}^{\star\star}$ satisfies the guarantee of Theorem 2.5. In the end we will use the union bound.
Suppose we have an estimator such that $\|\hat{\beta}_{j}-\beta\|_{2}\leq\tilde{d}_{j}$. Consider the set $K_{(j)}=K\cap B_{2}(0,\tilde{d}_{j}/2)$ and compute $m$ as in Corollary 5.1. Denote the corresponding matrix $A_{j}^{\star\star}=(I-X_{j}^{\star\star})^{1/2}$. Let $X_{j}^{\star\star}=V\Lambda V^{\top}$ be the eigendecomposition of $X_{j}^{\star\star}$.
We consider the estimate $\beta^{\prime}_{j+1}$ defined in (5.1). We will argue that by Theorem 2.5.9 of Dadush (2012) (see also Theorem 1 in Lee et al. (2018)) that $\beta^{\prime}_{j+1}$ is computable. First, we note that the set $K_{(j)}$ admits a membership oracle since we can evaluate its Minkowski gauge. Furthermore, the convex map $\nu\mapsto\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})-Z_{i}^{\top}A^{\star\star}\nu)_{i\in[N]}\|_{2}$ is Lipschitz. We have
$|\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})-Z_{i}^{\top}A^{\star\star}\nu)_{i\in[N]}\|_{2}-\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})-Z_{i}^{\top}A^{\star\star}\nu^{\prime})_{i\in[N]}\|_{2}|$
$\leq\|(Z_{i}^{\top}A^{\star\star}(\nu-\nu^{\prime}))_{i\in[N]}\|_{2}$
$\leq\sqrt{\lambda_{\max}\bigg(A^{\star\star}\sum_{i\in[N]}Z_{i}Z_{i}^{\top}A^{\star\star}\bigg)}\|\nu-\nu^{\prime}\|_{2}$
$\leq\sqrt{\lambda_{\max}\bigg(\sum_{i\in[N]}Z_{i}Z_{i}^{\top}}\bigg)\|\nu-\nu^{\prime}\|_{2}$
We have by Theorem 5.39 of Vershynin (2010):
$\sqrt{\lambda_{\max}\bigg(\sum_{i\in[N]}\Sigma^{-1/2}Z_{i}Z_{i}^{\top}\Sigma^{-1/2}\bigg)}\leq\sqrt{N}+C\sqrt{n}+t,$
with probability at least $1-2\exp(-ct^{2})$. Recall that $\lambda_{\max}(\Sigma^{-1/2})=\lambda_{\min}(\Sigma)^{-1/2}=O(1)$. Setting $t=\sqrt{N}\tilde{d}_{j}$ we obtain that with probability at least $1-2\exp(-cN\tilde{d}_{j}^{2})$
$\sqrt{\lambda_{\max}\bigg(\sum_{i\in[N]}Z_{i}Z_{i}^{\top}\bigg)}\lesssim\sqrt{N}+C\sqrt{n}+\sqrt{N}\tilde{d}_{j},$ (5.2)
and thus the map $\nu\mapsto\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})-Z_{i}^{\top}A^{\star\star}\nu)_{i\in[N]}\|_{2}$ is Lipschitz with a constant $\lesssim(\sqrt{N}+c\sqrt{n}+\sqrt{N}\tilde{d}_{j})\asymp(\sqrt{N}+\sqrt{n})$, since $\tilde{d}_{j}=O(1)$ by assumption. Hence the above approximate minimization over $K_{(j)}$ is computable by Theorem 2.5.9 in Dadush (2012).
Let $K^{\prime}_{(j)}=\{A_{j}^{\star\star}\nu:\nu\in K_{(j)}\}$. Denote with $\nu_{j}:=(\beta-\hat{\beta}_{j})/2\in K_{(j)}$ by assumption. Let $\Pi_{K^{\prime}_{(j)}}\nu_{j}=\mathop{\mathrm{argmin}}_{\nu\in K^{\prime}_{(j)}}\|\nu-\nu_{j}\|_{2}$. We know that $\|\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j}\|_{2}\leq\|A_{j}^{\star\star}\nu_{j}-\nu_{j}\|_{2}\lesssim\tilde{d}_{j}/c$ where the last inequality follows from a similar argument as in the proof of Lemma 2.9. Next using Theorem 2.5.9 Dadush (2012) with $\epsilon\asymp\sqrt{N}\tilde{d}_{j}^{2}$ we can obtain a near opitmal solution $\beta^{\prime}_{j+1}$. Then by the (near) optimality we have:
$\sum_{i\in[N]}((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})/2-Z_{i}^{\top}A_{j}^{\star\star}\beta^{\prime}_{j+1})^{2}$ $\leq\sum_{i\in[N]}((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})/2-Z_{i}^{\top}\Pi_{K^{\prime}_{(j)}}\nu_{j})^{2}+\epsilon^{2}$
$+2\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})/2-Z_{i}^{\top}\Pi_{K^{\prime}_{(j)}}\nu_{j})_{i\in[N]}\|_{2}\epsilon.$ (5.3)
Next using the shorthand $\hat{\Sigma}:=N^{-1}\sum_{i\in[N]}Z_{i}Z_{i}^{\top}$, we handle the last term of the right hand side:
$\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})/2-Z_{i}^{\top}\Pi_{K^{\prime}_{(j)}}\nu_{j})_{i\in[N]}\|_{2}\leq 2\sqrt{N(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})^{\top}\hat{\Sigma}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})}+2\|\xi\|_{2}.$
Using (5.2), the first term can be bounded as $\sqrt{N(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})^{\top}\hat{\Sigma}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})}\lesssim(\sqrt{N}+\sqrt{n})\tilde{d}_{j}/c$. On the other hand, by Proposition 5.16 of Vershynin (2010) we know that
$\mathbb{P}(\|\xi\|^{2}_{2}-\mathbb{E}\|\xi\|^{2}_{2}\geq t)\leq 2\exp(-C^{\prime\prime}(t^{2}/N\wedge t)).$
Hence we can select $t=N\tilde{d}_{j}$ to obtain that with probability at least $1-2\exp(-C^{\prime\prime}N\tilde{d}_{j}^{2})$ we have $\|\xi\|^{2}_{2}\leq\mathbb{E}\|\xi\|_{2}^{2}+N\tilde{d}_{j}\leq N\gamma^{2}+N\tilde{d}_{j}\lesssim N$. Hence since $\epsilon\asymp\sqrt{N}\tilde{d}_{j}^{2}$, and $\tilde{d}_{j}=O(1)$ we have that:
$\epsilon^{2}+2\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})/2-Z_{i}^{\top}\Pi_{K^{\prime}_{(j)}}\nu_{j})_{i\in[N]}\|_{2}\epsilon=O(N\tilde{d}_{j}^{2}).$
with high probability.
Then by rearranging (5) we have
$(\tilde{\beta}_{j+1}-\nu_{j})^{\top}\hat{\Sigma}(\tilde{\beta}_{j+1}-\nu_{j})$
$\leq(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})^{\top}\hat{\Sigma}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})+2\bigg|N^{-1}\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\tilde{\beta}_{j+1})\bigg|+\tilde{d}_{j}^{2}/C^{{}^{\prime}2},$
where $\tilde{\beta}_{j+1}=A_{j}^{\star\star}\beta^{\prime}_{j+1}$. Since $\sqrt{a+b+c}\leq\sqrt{a}+\sqrt{b}+\sqrt{c}$ for any $a,b,c\geq 0$ we have
$\|\hat{\Sigma}^{1/2}(\tilde{\beta}_{j+1}-\nu_{j})\|_{2}\leq\|\hat{\Sigma}^{1/2}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})\|_{2}+\sqrt{\bigg|2N^{-1}\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\tilde{\beta}_{j+1})\bigg|}+\tilde{d}_{j}/C^{\prime}.$ (5.4)
We first control the second term on the right hand side. Let $\tilde{\nu}_{j}\in K_{(j)}$, so that $\Pi_{K^{\prime}_{(j)}}\nu_{j}=A^{\star\star}_{j}\tilde{\nu}_{j}$. Also, let $A_{j}^{\star\star}=\sum_{k\in[n]}\sqrt{1-\lambda_{k}}v_{k}v_{k}^{\top}$ be the eigendecomposition of $A_{j}^{\star\star}$.
We can now bound the second term as follows:
$\bigg|\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\tilde{\beta}_{j+1})\bigg|$ $=\bigg|\sum_{i\in[N]}\sum_{k\in[n]}\xi_{i}Z_{i}^{\top}\sqrt{1-\lambda_{k}}v_{k}v_{k}^{\top}(\tilde{\nu}_{j}-\beta^{\prime}_{j+1})\bigg|$
$\leq\sqrt{\sum_{k}(1-\lambda_{k})\cdot(\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}v_{k})^{2}}\|\tilde{\nu}_{j}-\beta^{\prime}_{j+1}\|_{2}$
$\leq\sqrt{\sum_{k}(1-\lambda_{k})\cdot\max_{k\in[n]}(\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}v_{k})^{2}}\|\tilde{\nu}_{j}-\beta^{\prime}_{j+1}\|_{2}$
$\leq\sqrt{\lceil N\tilde{d}_{j}^{2}/C^{2}\rceil\cdot\max_{k\in[n]}(\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}v_{k})^{2}}\underbrace{(\tilde{d}_{j}/2\wedge\mathrm{diam}(K))}_{O(1)}$
$\lesssim\sqrt{N\tilde{d}_{j}^{2}/C^{2}\cdot\max_{k\in[n]}(\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}v_{k})^{2}}\underbrace{(\tilde{d}_{j}/2\wedge\mathrm{diam}(K))}_{O(1)},$ (5.5)
where in the last inequality we used that $\tilde{d}_{j}\geq C/\sqrt{N}$. Since both $\xi_{i}$ and $Z_{i}^{\top}v_{k}$ are sub-Gaussian variables their product is sub-exponential, and the constant is $O(1)$ by assumption. Moreover, clearly, $\mathbb{E}\xi_{i}Z_{i}^{\top}v_{k}=0$. Thus
$\mathbb{P}\bigg(\max_{k\in[n]}N^{-1/2}\bigg|\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}v_{k}\bigg|\geq t\bigg)\leq n\exp(-L\cdot(t^{2}\wedge\sqrt{N}t)),$
by Proposition 5.16 of Vershynin (2010). Set $t=\sqrt{N}\tilde{d}_{j}$, and note that $\tilde{d}_{j}=O(1)$ thus:
$\mathbb{P}\bigg(\max_{k\in[n]}N^{-1/2}\bigg|\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}v_{k}\bigg|\geq\sqrt{N}\tilde{d}_{j}\bigg)\leq n\exp(-L^{\prime}\cdot N\tilde{d}_{j}^{2}).$
Thus, continuing bound (5), we conclude that
$\bigg|\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\tilde{\beta}_{j+1})\bigg|\leq N\tilde{d}_{j}^{2}/C$
with probability at least $1-n\exp(-L^{\prime}\cdot N\tilde{d}_{j}^{2})$.
Once again by Theorem 5.39 of Vershynin (2010), we have that
$1-C\frac{\sqrt{n}}{\sqrt{N}}-\frac{t}{\sqrt{N}}\leq\sqrt{\lambda_{\min}(\Sigma^{-1/2}\hat{\Sigma}\Sigma^{-1/2})}\leq\sqrt{\lambda_{\max}(\Sigma^{-1/2}\hat{\Sigma}\Sigma^{-1/2})}\leq 1+C\frac{\sqrt{n}}{\sqrt{N}}+\frac{t}{\sqrt{N}},$
with probability at least $1-2\exp(-ct^{2})$. Selcting $t=\sqrt{N}\tilde{d}_{j}/C^{\prime}$ for some large $C^{\prime}$ we obtain
$1\asymp\lambda_{\min}(\Sigma^{-1/2}\hat{\Sigma}\Sigma^{-1/2})\asymp\lambda_{\max}(\Sigma^{-1/2}\hat{\Sigma}\Sigma^{-1/2}),$
under our assumption $N\gtrsim n$ with probability at least $1-\exp(-c^{\prime}N\tilde{d}_{j}^{2})$.
Combining, our previous inequalities with (5.4), we can claim that
$\lambda_{\min}(\Sigma)^{1/2}\|\tilde{\beta}_{j+1}-\nu_{j}\|_{2}$ $\lesssim\lambda_{\max}(\Sigma)^{1/2}\|\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j}\|_{2}+d_{j}/T^{\prime}+d_{j}/C^{\prime}$
$\leq d_{j}/T$
for some constants $T$ and $T^{\prime}$ that depend on $C,\lambda_{\max}(\Sigma)=O(1),c,n/N$ and other constants defined in this proof. Thus since $\lambda_{\min}(\Sigma)^{1/2}=\Omega(1)$ we have
$\|\tilde{\beta}_{j+1}-\nu_{j}\|_{2}\leq d_{j}/L,$
for some constant $L$ which can be made arbitrarily small by making the constant $C$ large. Upon grouping all of our probability bounds we obtain that this happens with probability at least $1-n\exp(-c^{\prime}N\tilde{d}_{j}^{2})$ for some constant $c^{\prime}$. The union bound completes the proof. ∎
Now that we have established Lemma 5.2, all that remains is to state the algorithm for estimating $\beta$. The algorithm is very similar to the one in the mean estimation setting:
[…]
5 Near-optimal Regression Estimation Algorithm
0: Convex body $K$, $r,R\in\mathbb{R}_{+}$ such that $rB_{2}\subseteq K\subseteq RB_{2}$, with $R=O(1)$; observations $(Y_{i},Z_{i})_{i\in[N]}$, constant $\tilde{L}=L/(\sqrt{3}+1)$ where $L$ is from Lemma 5.2; We assume $\tilde{L}>2(\sqrt{3}+1)$
0: Estimator $\hat{\beta}$ of $\beta$
1: Set $r=r\wedge(1/2\sqrt{n/N})$
2: Set $\tilde{d}_{1}=2R$, $K_{(1)}:=K$, number of iterations $M=\lceil\log_{\tilde{L}/(2(\sqrt{3}+1))}R/r\rceil$
3: Set $\hat{\beta}_{1}:=0$
4:for j = 1, …, M do
5:  Find $m=\lceil N\tilde{d}_{j}^{2}/C^{2}\rceil\wedge n$, and calculate the matrix $A_{j}^{\star\star}$ for the set $K_{(j)}$
6:  Find $\tilde{\beta}_{j+1}$, the estimator from Lemma 5.2
7:  Set $\bar{\beta}_{j+1}=\Pi_{K_{(j)}}\tilde{\beta}_{j+1}$ [^18] Here we set $\epsilon=\tilde{d}_{j}/L$ in Lemma A.1 when evaluating the weak projection.
8:  Lift back to $K$: $\hat{\beta}_{j+1}=\Pi_{K}(2\bar{\beta}_{j+1}+\hat{\beta}_{j})$ [^19] Here we set $\epsilon=2\tilde{d}_{j}/\tilde{L}$ in Lemma A.1 when evaluating the weak projection.
9:  Set $\tilde{d}_{j+1}=2(\sqrt{3}+1)\tilde{d}_{j}/\tilde{L}$
10:  Set $K_{(j+1)}:=K\cap B_{2}(0,\tilde{d}_{j+1}/2)$
11:  if $\tilde{d}_{j+1}\leq(2r)\vee(C/\sqrt{N})$ then
12:   return $\hat{\beta}:=\hat{\beta}_{j+1}$
13:  end if
14:end for
15:return $\hat{\beta}:=\hat{\beta}_{M+1}$
Since the justification of Algorithm 5 is very similar to the proof of Theorem 2.11, we omit it. However we will highlight a crucial modification that is required. We note that by a result of Prasadan and Neykov (2025a) for ellipsoidal constraints in linear models (see Section 5.1.2 therein) we have that the minimax rate on the squared level ($\inf_{\hat{\beta}}\sup_{\beta\in B_{2}(0,\tilde{d}_{k}/2)}\mathbb{E}\|\hat{\beta}-\beta\|_{2}^{2}$) for a ball $B_{2}(0,\tilde{d}_{k}/2)$ is of the order of $n/N\wedge\tilde{d}^{2}_{k}/4$. Now, since $r<\sqrt{n}/\sqrt{N}$ (here we mean the overwritten value of $r$) and $\tilde{d}_{k}\leq 2r$ then the minimax rate reduces to $c^{\prime}\tilde{d}_{k}^{2}$, for some $c^{\prime}>0$.
We can now state the main theorem of this section.
Theorem 5.3.
Suppose that $T_{2}(K)$ and $\max_{i\in[M]}\kappa(K_{(k)})$ are “small” factors and that $\log R/r$ is polynomial in $n,N$. Let $N\gtrsim n$ for an absolute constant. Then there exists a polynomial up to logarithmic factors function $\Upsilon(K)=g(T_{2}(K),\max_{i\in[M]}\kappa(K_{(k)}),\log n)$ [^20] Here we have $\log n$ in the function $g$ since the high probability in Lemma 5.2 contains an $n$ in front of the exponential which will result in a loss of logarithmic factor in $n$, i.e., we will have a term $\sqrt{(\log n)/N}$ added to the rate. We can now use the fact that $\eta^{\star}(K)\gtrsim 1/\sqrt{N}\wedge\mathrm{diam}(K)$. One can prove this similarly to the proof of Lemma 1.4 of Prasadan and Neykov (2025b). In the case when $\mathrm{diam}(K)\lesssim 1/\sqrt{N}$ it is obvious that our estimator is minimax optimal. such that the output $\hat{\beta}$ of Algorithm 2 satisfies
$\mathbb{E}\|\hat{\beta}-\beta\|_{2}^{2}\lesssim\Upsilon(K)\eta^{\star 2}(K),$
and is thus nearly minimax optimal.
Since ellipsoids are type-2 sets with small type-2 constant, we note that our method provides efficient implementation for linear regression problem under ellipsoidal constraints considered in the recent paper Pathak et al. (2024), provided that our further assumptions on the constraint set, covariates and noise vector are satisfied.
6 Some Remarks
In this section, we give an example of a type of convex bodies that obey all conditions previously required. It has long been known Donoho et al. (1990) that for convex Quadratically Convex Orthosymmetric (QCO) sets, the minimax rate is given by an equation involving the Kolmogorov widths. In fact, the truncated series estimator is rate optimal over any QCO set. However, it was not previously known how one can compute a minimax optimal estimator (including the truncated series one) for QCO sets in generality. Here we argue that any orthosynmetric QCO set induces a norm which is of type-2 with small type-2 constant and is exactly 2-convex.
Lemma 6.1.
Let $K\subset\mathbb{R}^{n}$ be a quadratically convex orthosymmetric (QCO) set, i.e.
  1. 1.
    $K$ is convex, origin-symmetric, and sign-invariant;
  2. 2.
    the set $K^{(2)}:=\{x^{2}:x\in K\}$ is convex, where squaring is done coordinatewise.
Then the norm $\rho_{K}(\cdot)$ induced by $K$ satisfies a type-2 inequality with constant $C\log n$:
$\mathbb{E}\rho_{K}(\sum_{i=1}^{m}\varepsilon_{i}x_{i})^{2}\;\leq\;C(\log n)\sum_{i=1}^{m}\rho_{K}(x_{i})^{2},$
for all $x_{1},\dots,x_{m}\in\mathbb{R}^{n}$, where $\varepsilon_{i}$ are i.i.d. Rademacher signs.
We note furthermore that the Minkowksi gauges of QCO sets are necessarily 2-convex. To show this define
$K^{2}:=\{\operatorname{sign}(x)\circ x^{2}:x\in K\}.$
Since $K$ is sign-invariant and quadratically convex, $K^{2}$ is convex and sign-invariant. By direct inspection,
$|x|\in tK\iff x^{2}\in t^{2}K^{2},$
and therefore
$\rho_{K}(x)=\sqrt{\rho_{K^{2}}(x^{2})}.$
Thus, by the triangle inequality:
$\rho_{K}^{2}((\sum_{i\in[m]}|x_{i}|^{2})^{1/2})=\rho_{K^{2}}(\sum_{i\in[m]}|x_{i}|^{2})\leq\sum_{i\in[m]}\rho_{K^{2}}(|x_{i}|^{2})=\sum_{i\in[m]}\rho_{K}^{2}(x_{i})$
It follows immediately that if $K\subset\mathbb{R}^{m}$ is a QCO set (with $m$ being “small”), we can implement our algorithm over bounded, well-balanced sets $K^{\prime}=\{x\in\mathbb{R}^{n}:\rho_{K}(Ax)\leq 1\}$ for some full rank matrix $A\in\mathbb{R}^{m\times n}$ (provided that we can evaluate $\rho_{K}$). Clearly in the special case when $A=I_{n}$, this implies that we can implement our algorithm for any well-balanced QCO set with oracle evaluating its Minkowski gauge. Examples of QCO include ellipsoids, hyperrectangles and sets of the type $\{x\in\mathbb{R}^{n}:\sum_{i\in[n]}\theta_{i}\psi(x_{i}^{2})\leq 1\}$ for some convex function $\psi$ and constants $\theta_{i}\in\mathbb{R}_{+}$.
It is easy to see that not all bodies with small type-2 constant are QCO sets. We attach the following proposition which verifies this fact.
Proposition 6.2.
Let $n\geq 2$ and define, for $x\in\mathbb{R}^{n}$,
$\|x\|=\Big(\|x\|_{2}^{2}+\frac{1}{n}\|x\|_{1}^{2}\Big)^{1/2}.$
Then $\|\cdot\|$ is a norm which is convex, sign–symmetric, permutation–invariant, and has type–2 constant bounded by an absolute constant. However, its unit ball
$K=\bigl\{x\in\mathbb{R}^{n}:\|x\|\leq 1\bigr\}$
is not quadratically convex.
Hence our algorithm can be implemented for well-balanced sets $K^{\prime}=\{x\in\mathbb{R}^{n}:\|Ax\|\leq 1\}$ where $A\in\mathbb{R}^{n\times n}$ is a full rank matrix. We conclude that our algorithm applies not only to transformed QCO sets, but also to a broader class of constraint sets.
7 Discussion
This paper proposed the first computationally efficient algorithm for constrained mean estimation in a Gaussian sequence model setting, heavy-tailed robust mean estimation and linear regression under type-2 convex constraints, where the constraint set satisfies further regularity assumptions as outlined in Sections 2 and 3. There is one central open question that remains: can one develop near optimal polynomial-time algorithms for other (convex) constraints. Notably our algorithm does not work for $\ell_{p}$ balls for $1\leq p<2$, for which efficient near-minimax optimal soft and hard thresholding algorithms do exist Johnstone (2019). We further had to assume that the set $K$ is origin-symmetric. The hypothetical efficient algorithm will likely have to use a fundamentally different approach as we are no longer guaranteed that the Kolmogorov widths decay as in the origin-symmetric type-2 sets considered in this paper.
Another interesting open question is whether our algorithm is meaningful for misspecified scenarios, i.e., when $\mu$ or $\beta$ do not belong to the set $K$.
Yet another question is as follows. In the work of Neykov (2022) when the constraint set is convex, one can show that the estimator is adaptive to the location of the true mean $\mu\in K$. However, we suspect that the estimator provided in this paper is not adaptive. It will be interesting to look into this rigorously: either establish that the estimator proposed here is adaptive, or refute it and provide an adaptive estimator.
Another avenue for future research is to try to remove the condition $d^{2}/\sigma^{2}\lesssim\exp(N/C_{0})/N$ in the robust algorithm. We believe this assumption is not necessary, but we will likely need to develop a new algorithm which converges without the need of the pilot estimator based on the work of Depersin and Lecué (2022).
In the regression setting we had to assume the dimension is smaller than the sample size up to an absolute constant. It will be interesting if one can remove this condition. We crucially used in the proof of Lemma 5.2, to show that the eigenvalues of the sample covariance matrix are upper and lower bounded, and we do not know presently how to remove it.
Finally, our mean estimation procedures require knowledge of $\sigma$, or an accurate upper bound. Whether one can develop near optimal adaptive computationally tractable algorithms is an open problem.
8 Acknowledgments
The author is grateful to Sourav Chatterjee for encouraging him to think deeply about this problem. He also benefited from valuable discussions with Andrea Montanari, Siva Balakrishnan, and Ilias Diakonikolas, as well as from exploratory conversations with ChatGPT. Finally, the author would like to thank Vijay Bhattiprolu for a helpful email exchange, where Vijay clarified several points of the central paper to our work Bhattiprolu et al. (2021). Furthermore, the author thanks Daniel Dadush for clarifying a point in Theorem 2.5.9 in his PhD thesis.
Appendix A Supplemental Proofs
Proof of Lemma 2.3.
We form the Lagrangian
$\mathcal{L}(w,\theta,\alpha,\beta)=\frac{1}{2}\|w-v\|_{2}^{2}+\theta\Big(\sum_{i=1}^{n}w_{i}-k\Big)-\sum_{i=1}^{n}\alpha_{i}w_{i}+\sum_{i=1}^{n}\beta_{i}(w_{i}-1),$
where $\alpha_{i},\beta_{i}\geq 0$ are the multipliers for the box constraints $w_{i}\geq 0$ and $w_{i}\leq 1$, respectively.
Stationarity gives
$0=\nabla_{w}\mathcal{L}=w-v+\theta\mathbf{1}-\alpha+\beta,\quad\text{i.e.}\quad w_{i}=v_{i}-\theta+\alpha_{i}-\beta_{i}.$
The complementary slackness conditions are:
$\alpha_{i}w_{i}=0,\qquad\beta_{i}(w_{i}-1)=0.$
Hence, for each coordinate $i$:
  • If $0<w_{i}<1$, then $\alpha_{i}=\beta_{i}=0$ and $w_{i}=v_{i}-\theta$.
  • If $w_{i}=0$, then $\alpha_{i}\geq 0$ and $v_{i}-\theta\leq 0$.
  • If $w_{i}=1$, then $\beta_{i}\geq 0$ and $v_{i}-\theta\geq 1$.
Therefore,
$w_{i}=\min\{1,\max\{0,v_{i}-\theta^{\star}\}\}.$
The equality constraint $\sum_{i}w_{i}=k$ then becomes
$f(\theta)=\sum_{i=1}^{n}\min\{1,\max\{0,v_{i}-\theta\}\}=k.$
The function $f(\theta)$ is continuous, strictly decreasing, and piecewise linear. As $f(\min(v)-1)=n$ and $f(\max(v))=0$, by the intermediate value theorem there exists a unique $\theta^{\star}$ such that $f(\theta^{\star})=k$. Substituting this $\theta^{\star}$ yields the claimed projection. Convexity of both the objective and constraint set ensures uniqueness. The last implication is obvious due to the formula for $w_{i}$. ∎
Lemma A.1.
Given a convex body with a weak separation oracle, we can weakly project on $K$, in the sense that for any $z\in\mathbb{R}^{n}$, and $\epsilon>0$ we can find a rational number $w$ and a point $p\in K$ in polynomial-time, such that
$w-\epsilon\leq\|z-\overline{\Pi}_{K}(z)\|_{2}\leq\|z-p\|_{2}\leq w,$
where $\overline{\Pi}_{K}(z)$ denotes the true Euclidean projection of $z$ on $K$. Moreover, for any point $\nu\in K$ we have
$\|p-\nu\|_{2}\leq\|z-\nu\|_{2}+\sqrt{\epsilon^{2}+2\epsilon\|z-\nu\|}.$
Corollary A.2.
Hence if $\|z-\nu\|\leq L$ and $\epsilon\leq L$ we have
$\|p-\nu\|_{2}\leq(\sqrt{3}+1)L$
Proof.
This is a direct corollary of Theorem 2.5.9 of [5], see also [17] using the fact that the $\ell_{2}$ norm is $1$-Lipschitz.
For the second part we first observe that
$\|\nu-\Pi_{K}(z)\|_{2}\leq\|z-\nu\|_{2}$
Next by the properties of the Euclidean projection on convex bodies we have:
$\|p-\Pi_{K}(z)\|_{2}^{2}+\|z-\Pi_{K}(z)\|_{2}^{2}\leq\|z-p\|_{2}^{2}\leq\|z-\Pi_{K}(z)\|_{2}^{2}+\epsilon^{2}+2\epsilon\|z-\Pi_{K}(z)\|_{2}$
Thus
$\|p-\Pi_{K}(z)\|_{2}\leq\sqrt{\epsilon^{2}+2\epsilon\|z-\Pi_{K}(z)\|_{2}}\leq\sqrt{\epsilon^{2}+2\epsilon\|z-\nu\|_{2}}\leq\sqrt{3}L,$
where we used the inequlity $\|z-\nu\|_{2}\vee\epsilon\leq L$. By the triangle inequality
$\|\nu-p\|\leq\|\nu-\Pi_{K}(z)\|_{2}+\|p-\Pi_{K}(z)\|_{2}\leq\|z-\nu\|_{2}+\sqrt{3}L\leq(\sqrt{3}+1)L.$
Proof of Lemma 6.1.
Step 1: Quadratic representation of the norm. Let $\|\cdot\|_{K}=\rho_{K}(\cdot)$. Define
$K^{2}:=\{\operatorname{sign}(x)\circ x^{2}:x\in K\}.$
Since $K$ is sign-invariant and quadratically convex, $K^{2}$ is convex, sign-invariant, and symmetric. By direct inspection,
$|x|\in tK\iff x^{2}\in t^{2}K^{2},$
and therefore
$\|x\|_{K}=\sqrt{\|x^{2}\|_{K^{2}}}.$ (A.1)
Step 2: Reduction to Gaussian averages. By [16, Lemma 4.5],
$\mathbb{E}\Big\|\sum_{i=1}^{m}\varepsilon_{i}x_{i}\Big\|_{K}^{2}\leq\frac{\pi}{2}\,\mathbb{E}\Big\|\sum_{i=1}^{m}g_{i}x_{i}\Big\|_{K}^{2},$
where $g_{i}\sim N(0,1)$ are independent Gaussians.
Step 3: Coordinatewise representation. Let $S:=\sum_{i=1}^{m}g_{i}x_{i}\in\mathbb{R}^{n}$. Then for each coordinate $j$,
$S_{j}=\sum_{i=1}^{m}g_{i}x_{ij}\sim N(0,\sigma_{j}^{2}),\quad\sigma_{j}^{2}=\sum_{i}x_{ij}^{2}.$
Hence $S_{j}^{2}=\sigma_{j}^{2}\chi_{j}^{2}$, where $\chi_{j}^{2}\sim\chi^{2}_{1}$, and in vector form
$S^{2}=\chi^{2}\circ\Big(\sum_{i}x_{i}^{2}\Big),$
where $\circ$” denotes coordinatewise multiplication.
Step 4: Monotonicity of sign-invariant norms. For any unconditional norm,
$|u|\leq|v|\implies\|u\|_{K^{2}}\leq\|v\|_{K^{2}}.$
Therefore,
$\|S^{2}\|_{K^{2}}\leq\big(\max_{j}\chi_{j}^{2}\big)\,\Big\|\sum_{i}x_{i}^{2}\Big\|_{K^{2}}.$
Step 5: Expectation bound. Taking expectations and using subexponential tails of $\chi^{2}_{1}$:
$\mathbb{E}\|S^{2}\|_{K^{2}}\leq\mathbb{E}\max_{j}\chi_{j}^{2}\cdot\Big\|\sum_{i}x_{i}^{2}\Big\|_{K^{2}},$
and
$\mathbb{E}\max_{1\leq j\leq n}\chi_{j}^{2}\lesssim\log n.$
Hence
$\mathbb{E}\|S^{2}\|_{K^{2}}\lesssim(\log n)\Big\|\sum_{i}x_{i}^{2}\Big\|_{K^{2}}.$ (A.2)
Step 6: Subadditivity and conclusion. By convexity of the norm,
$\Big\|\sum_{i}x_{i}^{2}\Big\|_{K^{2}}\leq\sum_{i}\|x_{i}^{2}\|_{K^{2}}.$
Using (A.1) and (A.2),
$\mathbb{E}\|S\|_{K}^{2}=\mathbb{E}\|S^{2}\|_{K^{2}}\lesssim(\log n)\sum_{i}\|x_{i}\|_{K}^{2}.$
Combining with step 2 yields the desired type-2 inequality. ∎
Proof of Proposition 6.2.
For $u=(1,0,\dots,0)$ and $v=(0,1,0,\dots,0)$, we have
$\|u\|_{2}=\|v\|_{2}=1,\qquad\|u\|_{1}=\|v\|_{1}=1,$
hence
$\|u\|=\|v\|=\sqrt{1+\frac{1}{n}}=:c_{n}>1.$
Let $\tilde{u}=u/c_{n}$ and $\tilde{v}=v/c_{n}$. Then $\|\tilde{u}\|=\|\tilde{v}\|=1$, so $\tilde{u},\tilde{v}\in K$.
Consider the midpoint in squared coordinates for $t=\tfrac{1}{2}$:
$\frac{1}{2}(\tilde{u}^{2}+\tilde{v}^{2})=\frac{1}{2c_{n}^{2}}(1,1,0,\dots,0),$
and take the entrywise square root
$w=\sqrt{\tfrac{1}{2}(\tilde{u}^{2}+\tilde{v}^{2})}=\frac{1}{c_{n}\sqrt{2}}(1,1,0,\dots,0).$
We now compute $\|w\|$:
$\|w\|_{2}^{2}=\frac{1}{c_{n}^{2}},\qquad\|w\|_{1}^{2}=\Big(\frac{2}{c_{n}\sqrt{2}}\Big)^{2}=\frac{2}{c_{n}^{2}}.$
Therefore
$\|w\|^{2}=\|w\|_{2}^{2}+\frac{1}{n}\|w\|_{1}^{2}=\frac{1}{c_{n}^{2}}\Big(1+\frac{2}{n}\Big)=\frac{1+\tfrac{2}{n}}{1+\tfrac{1}{n}}=\frac{n+2}{n+1}=1+\frac{1}{n+1}.$
Hence $\|w\|=\sqrt{1+\tfrac{1}{n+1}}>1$, so $w\notin K$.
Thus, although $\tilde{u},\tilde{v}\in K$, the vector $\sqrt{(\tilde{u}^{2}+\tilde{v}^{2})/2}$ leaves the unit ball. The latter shows that $K$ is not quadratically convex.
Finally, since
$\|x\|_{2}\leq\|x\|\leq\sqrt{2}\|x\|_{2},$
the norm is O(1)O(1)–equivalent to the Euclidean norm and therefore has type–2 constant bounded by an absolute constant. ∎